### v1

# Moment Tensor Potential and Equivariant Tensor Network Potential with explicit dispersion interactions

Abstract
In this study, we investigate the effect of incorporating explicit dispersion interactions in the functional form of machine learning interatomic potentials (MLIPs), particularly in the Moment Tensor Potential and Equivariant Tensor Network potential for accurate modeling of liquid carbon tetrachloride, methane, and toluene. We show that explicit incorporation of dispersion interactions via D2 and D3 corrections significantly improves the accuracy of MLIPs when the cutoff radius is set to a commonly used value of 5—6 Å. We also show that for carbon tetrachloride and methane, a substantial improvement in accuracy can be achieved by extending the cutoff radius to 7.5 Å. However, for accurate modeling of toluene, explicit incorporation of dispersion remains important. Furthermore, we find that MLIPs incorporating dispersion interactions via D2 reach a close level of accuracy to those incorporating D3, and D2 is suitable for accurate modeling of the systems in the study, while being less computationally expensive. We evaluated the accuracy of MLIPs in dimer binding curves compared to ab initio data and in predicting density and radial distribution functions compared to experiments.
preprint: AIP/123-QED
I Introduction
Machine learning interatomic potentials (MLIPs) have emerged as widely recognized tools for modeling large systems over an extended timescale with the accuracy of ab initio methods. MLIPs typically achieve relative errors in energy and forces of a few percent, evaluated with respect to ab initio results, while also demonstrating a favorable scaling of computational complexity as the number of atoms increasesZuo et al. (2020); Kozinsky et al. (2023); Zhang et al. (2025). This makes MLIPs an attractive choice for simulating systems that are, on the one hand, computationally prohibitive for conventional ab initio methods and, on the other hand, beyond the accuracy of (semi-)empirical force fields.
Despite significant advances in the development of MLIPs, several key challenges remain. The primary challenge for MLIPs arises from their locality, which is a consequence of the introduction of a cutoff radius that restricts the number of interacting atomsZhang et al. (2025). Although this approximation works well for metals and alloys, it requires greater caution for low-dimensional materials, materials with defects, or moleculesBabaei and Sadeghi (2024); Behler and Csányi (2021); Anstine and Isayev (2023). Moreover, the locality of MLIPs restricts them from accurately capturing long-range interactions, such as electrostatic, induction, and dispersionAnstine and Isayev (2023).
MLIPs that explicitly incorporate long-range interactions have already been presented in the literature. Such MLIPs include PhysNetUnke and Meuwly (2019), TensorMolYao et al. (2018), and ee4G-HDNNP model Ko et al. (2023), which incorporate both electrostatic and dispersion interactions. Other MLIPs such as HDNNP potentialsArtrith, Morawietz, and Behler (2011); Ko et al. (2021), and DPLR modelZhang et al. (2022) incorporate electrostatic interactions. At the same time, GAPBartók et al. (2010) was applied to study systems with strong dispersion interactionsVeit et al. (2019); Rowe et al. (2020, 2018); Ibragimova et al. (2025). However, it is also possible to model intramolecular and intermolecular interactions using separate potentialsWengert et al. (2021); Gao and Remsing (2022). Additionally, an alternative approach is to account for long-range interactions via global descriptorsKabylda et al. (2023).
The most common approach to address long-range interactions is to explicitly incorporate physically motivated terms in the total energy expressionAnstine and Isayev (2023). Behler proposed a classification of MLIPs into four generations based on how long-range interactions are incorporatedBehler (2021); Behler and Csányi (2021). The first and second generations correspond to purely local potentials, while the third generation MLIPs explicitly include long-range interactions, with their strength defined by local features such as atomic charges, dependent on the local environment. The fourth generation accounts for the nonlocality of parameters defining long-range interactions, particularly by making charges dependent on the nonlocal environment. This classification is widely used in the context of MLIPs that incorporate electrostatic interactions and can also be applied to classify MLIPs that account for dispersion interactionsBehler and Csányi (2021).
For explicit capture of dispersion interactions, the most common approach is to utilize one of the dispersion corrections developed for density functional theory (DFT) calculationsGrimme et al. (2016); Hermann, DiStasio Jr, and Tkatchenko (2017), such as Tkatchenko-Scheffler (TS) Tkatchenko and Scheffler (2009), exchange-dipole moment (XDM)Becke and Johnson (2005), D3Grimme et al. (2010) or D2Grimme (2006) models. For electronic density-dependent dispersion correction schemes, such as TS or XDM, machine learning can be used to predict the most computationally demanding parameters, such as atomic moments and volumes in XDMTu et al. (2023, 2024). Further optimization of computational cost can be achieved using averaged parameters instead of machine learning: atomic volumes for TSVeit et al. (2019); Muhli et al. (2021); Ibragimova et al. (2025) or moments and atomic volumes for XDMTu et al. (2023). The latter approach has been shown to have accuracy close to the one of the scheme using machine-learned coefficientsTu et al. (2023). The D2 or D3 dispersion corrections, which do not require electronic density for evaluation, can be applied on top of a local potentialWang and Huang (2020); Ying and Fan (2023); Unke and Meuwly (2019); Yao et al. (2018), without machine learning of the dispersion correction parameters. An alternative involves parameterizing the Lennard-Jones potential based on relevant configurations, such as the exfoliation curve for black phosphorus or grapheneRowe et al. (2018); Deringer, Caro, and Csányi (2020).
Another challenge is the efficient generation of high-quality ab initio training sets, which is crucial to ensure accuracy within the training domain, for efficient construction of training set, and to improve the transferability of the resulting potential. To address this challenge, various data-driven approaches have been developed, focusing on optimization of training sets rather than increasing model complexity. These approaches include different variations of active learning (AL), such as methods based on the Query by Committee strategySmith et al. (2018), Bayesian inferenceVandermause et al. (2020), optimization of entropy of atomic descriptorsMontes de Oca Zapiain et al. (2022), and the D-optimality criterion employed together with the MaxVol algorithmPodryabinkin and Shapeev (2016).
In the present work, we focus on the development of MLIPs with explicit treatment of the dispersion interaction, particularly Moment Tensor Potentials (MTPs)Shapeev (2016) and Equivariant Tensor Network (ETN) potentialsHodapp and Shapeev (2024). MTP is implemented in the MLIP-2 software, together with the AL approach based on the D-optimality criterionNovikov et al. (2020), which enables efficient training set generation for MTP. ETN is implemented in the MLIP-4 softwareShapeev (2024). MTP has been widely applied to the modeling of inorganic materialsKorotaev et al. (2019); Podryabinkin et al. (2019); Wang et al. (2020), metals and alloysMarmolejo-Tejada and Mosquera (2022); Tasnadi et al. (2021); Gil and Oda (2022); Zheng et al. (2023), and molten saltsRybin et al. (2024); Sun et al. (2024); Attarian, Morgan, and Szlufarska (2022). However, it was not extensively used to study molecular compoundsRybin, Novikov, and Shapeev (2025); Klimanova, Rybin, and Shapeev (2024); Reicht et al. (2024); Novikov et al. (2024). At the same time, ETN has recently been developedHodapp and Shapeev (2024) and has been successfully applied to defects in metallic alloysHodapp (2025), but has not yet been widely benchmarked in various applications.
Here, our objective is to improve the accuracy and applicability of the MTP and ETN models by explicitly incorporating dispersion interactions via D2 and D3 corrections. In our study, we compare the precision of local MLIPS (MLIPS(l)) with MLIPS with explicit dispersion correction (MLIP+D). We study the accuracy differences between MLIP(l)’s and MLIP+D’s when the cutoff radius is set to a value within a commonly used range of 5—6 Å, thereby constraining the MLIP(l)’s ability to capture the full range of dispersion interactions. We additionally test an alternative approach in which we simply extend the cutoff radius for MLIP(l)’s to 7.5 Å, allowing them to describe the nearly full range of dispersion interactions, and compare the accuracy of the resulting MLIP(l)’s with MLIP+D’s.
This comparative analysis was performed on liquid carbon tetrachloride, methane, and toluene simulations, since these systems exhibit predominant dispersion interactions and are widely represented in the literature. Moreover, carbon tetrachloride and toluene are common nonpolar solvents. To examine the predictive power of potentials, we calculate binding curves and compare them with the ones obtained with ab initio calculations. We also calculate density and radial distribution function (RDF) with the fitted MLIPs and compare them to the experimental ones.
II Methodology
II.1 Moment Tensor Potential
Moment Tensor Potential (MTP) is a machine learning interatomic potential which was first proposed for single-component materialsShapeev (2016) and generalized to multi-component materialsGubaev, Podryabinkin, and Shapeev (2018). MTP energy $E^{\rm MTP}$ is the sum of contributions $V^{\rm MTP}(\mathfrak{\bm{n}}_{i})$ of atomic neighborhoods ${\bf\mathfrak{n}}_{i}$ for $N$ atoms
$E^{\rm MTP}=\sum\limits_{i=1}^{N}V^{\rm MTP}(\mathfrak{\bm{n}}_{i}).$ (1)
Each neighborhood is a tuple
$\mathfrak{n}_{i}=(\{r_{i1},z_{i},z_{1}\},\ldots,\{r_{ij},z_{i},z_{j}\},\ldots,\{r_{iN_{\rm nbh}},z_{i},z_{N_{\rm nbh}}\}),$
where $r_{ij}$ are relative atomic positions, $z_{i}$, $z_{j}$ are the types of central and neighboring atoms, and $N_{\rm nbh}$ is the number of atoms in neighborhood. Each contribution $V^{\rm MTP}(\mathfrak{\bm{n}}_{i})$ in the potential energy $E^{\rm MTP}$ expands through a set of basis functions
$V^{\rm MTP}({\bf\mathfrak{n}}_{i})=\sum\limits_{\alpha}\xi_{\alpha}B_{\alpha}({\mathfrak{\bm{n}}}_{i}),$ (2)
where $B_{\alpha}$ are the MTP basis functions, $\xi_{\alpha}$ are the linear parameters to be found. To define the functional form of the MTP basis functions we introduce the so-called moment tensor descriptors:
$M_{\mu,\nu}({\mathfrak{\bm{n}}}_{i})=\sum_{j=1}^{N_{\rm nbh}}f_{\mu}(|r_{ij}|,z_{i},z_{j})r_{ij}^{\otimes\nu}.$ (3)
The descriptor consists of the angular part $r_{ij}^{\otimes\nu}$ (the symbol $\otimes$” denotes the outer product of vectors and, thus, the angular part is the tensor of $\nu$-th order) and the radial part $f_{\mu}(|r_{ij}|,z_{i},z_{j})$ of the following form:
$\displaystyle f_{\mu}(|r_{ij}|,z_{i},z_{j})=\sum_{\beta}c^{(\beta)}_{\mu,z_{i},z_{j}}T^{(\beta)}(|r_{ij}|)(R_{\rm cut}-|r_{ij}|)^{2}.$ (4)
Here $\mu$ is the number of the radial function $f_{\mu}$, $c=\{c^{(\beta)}_{\mu,z_{i},z_{j}}\}$ are the radial parameters to be found, $T^{(\beta)}(|r_{ij}|)$ are polynomial functions, and the term $(R_{\rm cut}-|r_{ij}|)^{2}$ is introduced to ensure smoothness with respect to the atoms leaving and entering the sphere with the cutoff radius $R_{\rm cut}$.
By definition, the MTP basis function $B_{\alpha}$ is a contraction of one or more moment tensor descriptors, yielding a scalar. In order to construct the basis functions $B_{\alpha}$, and, thus, determine a particular functional form of MTP, we define the so-called level of moment tensor descriptor:
${\rm lev}M_{\mu,\nu}=2+4\mu+\nu.$ (5)
We also define the level of MTP basis function:
${\rm lev}B_{\alpha}=\rm{lev}\underbrace{\prod_{p=1}^{P}M_{\mu_{p},\nu_{p}}}_{scalar}=\sum\limits_{p=1}^{P}(2+4\mu_{p}+\nu_{p}).$ (6)
A set of MTP basis functions and, thus, a particular functional form of MTP depends on the maximum level, ${\rm lev_{\rm max}}$, which we also call the level of MTP. In the set of MTP basis functions, we include only those with ${\rm lev}B_{\alpha}\leq{\rm lev_{\rm max}}$.
II.2 Equivariant Tensor Network Potentials
Polynomial-based potentials, like MTPs, can be generally written as a contraction of a tensor $T$, containing the parameters of the potential, with feature vectors $v$ that contain the same radial and angular information as the moment tensor descriptors (3) (we will make this precise further below)
$V({\bf\mathfrak{n}}_{i})=T_{k_{1}\ldots k_{d}}v^{1}_{k_{1}}\ldots v^{d}_{k_{d}},$ (7)
where the dimension $d$ defines the body-order of the potential. However, Equation (7) is not convenient because the evaluation scales exponentially with $d\cdot n$, where $n$ is the number of features contained in each of the $v$’s. For MTPs, this problem is to some extent solved semi-empirically, e.g., by contracting the radial features before entering the moment tensor descriptors.
A way to explicitly control the growth of the potential’s complexity is to represent $T$ with a tensor network. Tensor networks are a way of representing high-dimensional tensors in a low-rank format by factorizing the full tensor into smaller tensors, up to the order of three. There are different formats, with tensor trains, hierarchical Tucker, or PEPS, arguably being among the most popular onesCichocki et al. ; Orús , but their contractions with vectors $v^{1}_{k_{1}}$, $v^{2}_{k_{2}}$, etc., can all be realized by sequences of contractions of up-to-order-three tensors
$u^{1}_{k_{1}^{\prime}}$ $=T^{1}_{k_{1}^{\prime}k_{1}}v^{1}_{k_{1}},$
$u^{2}_{k_{2}^{\prime}}$ $=T^{2}_{k_{2}^{\prime}k_{2}k_{1}^{\prime}}v^{2}_{k_{2}}u^{1}_{k_{1}^{\prime}},$
$u^{3}_{k_{3}^{\prime}}$ $=T^{3}_{k_{3}^{\prime}k_{3}k_{2}^{\prime}}v^{3}_{k_{3}}u^{2}_{k_{2}^{\prime}},$
$\ldots$.… .
The main difference between ETNs and conventional tensor networks is the implementation of symmetry constraints that render the ETN invariant under the actions of the corresponding symmetry group. In our case of interatomic potentials, we require that the energy per atom $V({\bf\mathfrak{n}}_{i})$ remains invariant under the actions of the group of rotations SO(3). [^1] To encode the full O(3) invariance (rotations and reflections) into ETNs, it suffices that $V({\bf\mathfrak{n}}_{i})$ is a real quantity (more details are provided in the original paperHodapp and Shapeev (2024), section 3.4.2) To that end, we require the feature vectors to be SO(3)-equivariant vectors, that rotate correspondingly with a basis change under SO(3). In the following, we consider a decomposition of each $v$ into an irreducible covariant representation of SO(3) using spherical harmonics. Thus, each index $k$ of a feature vector $v_{k}$ is a multi-index $k=(\ell mn)$, with $\ell=0,\ldots,L$ being the index of the subspace of the irreducible representation, $m\in\{-\ell,-\ell+1,\ldots,\ell\}$ being the dimension of the subspace, and $n=1,\ldots,N(\ell)$ is the number of radial channels corresponding to each $\ell$. We point out that we intentionally deviate from the ordering $n\ell m$, commonly used in quantum physics, that puts the index $n$ in front of $\ell$ and $m$. We have chosen this notation because each multi-index in our tensor network may depend on a different $\ell$. Therefore, since $m$ and $n$ are always assumed to depend on $\ell$, the ordering $\ell mn$ appears to be more comprehensible in our context.
With this definition of the feature vectors, a sufficient condition for $V({\bf\mathfrak{n}}_{i})$ being invariant under actions of SO(3) is that the tensors $T$ are equivariant maps of these covariant vectors. According to the Wigner-Eckhart Theorem, any $T$ with three multi-indices $\{(\ell_{i},m_{i},n_{i})\}_{i=1,\ldots,3}$, can be factorized to
$T_{(\ell_{1}m_{1}n_{1})(\ell_{2}m_{2}n_{2})(\ell_{3}m_{2}n_{3})}=\theta_{(\ell_{1}n_{1})(\ell_{2}n_{2})(\ell_{3}n_{3})}C_{(\ell_{1}m_{1})(\ell_{2}m_{2})(\ell_{3}m_{3})},$
where $\theta_{(\ell_{1}n_{1})(\ell_{2}n_{2})(\ell_{3}n_{3})}$ is the tensor of model coefficients, and $C_{(\ell_{1}m_{1})(\ell_{2}m_{2})(\ell_{3}m_{3})}$ is the Clebsch-Gordan coefficient that defines the symmetry group. As a tensor network, we use an (equivariant) tensor train representation Oseledets (2011) of atomic energies with equal feature vectors $(v^{1},\ldots,v^{d})=v$. An ETN potential in the tensor train format can then be written as follows
$V^{\rm ETN}({\bf\mathfrak{n}}_{i})=\left(T^{1}_{(\ell_{1}^{\prime}m_{1}^{\prime}n_{1}^{\prime})(\ell_{1}m_{1}n_{1})}v_{(\ell_{1}^{\prime}m_{1}^{\prime}n_{1}^{\prime})}({\bf\mathfrak{n}}_{i})\right)\\ \left(T^{2}_{(\ell_{1}m_{1}n_{1})(\ell_{2}^{\prime}m_{2}^{\prime}n_{2}^{\prime})(\ell_{2}m_{2}n_{2})}v_{(\ell_{2}^{\prime}m_{2}^{\prime}n_{2}^{\prime})}({\bf\mathfrak{n}}_{i})\right)\\ \ldots\left(T^{d}_{(\ell_{d-1}m_{d-1}n_{d-1})(\ell_{d}^{\prime}m_{d}^{\prime}n_{d}^{\prime})}v_{(\ell_{d}^{\prime}m_{d}^{\prime}n_{d}^{\prime})}({\bf\mathfrak{n}}_{i})\right)$ (8)
in which the channels $n_{1},n_{2},\ldots,n_{d}$ naturally emerge as ranks of the tensor network. As feature vectors, we use the ones that contract the radial features before entering the tensor trainHodapp and Shapeev (2024) as follows:
$v_{(\ell mn)}=\sum_{j}\Big{(}B_{\ell n\alpha\lambda}Q_{\alpha}(|r_{ij}|)\big{(}A_{\ell\lambda\beta\gamma}z_{\beta}^{i}z_{\gamma}^{j}\big{)}\Big{)}Y_{\ell m}(r_{ij}/|r_{ij}|).$
When comparing the moment tensor descriptors $M_{\mu,\nu}({\mathfrak{\bm{n}}}_{i})$ and the feature vectors of ETNs, the radial parameters $c$ are decomposed into $A$ and $B$, and the angular part $r_{ij}^{\otimes\nu}$ is reshaped to $Y_{\ell m}(r_{ij}/|r_{ij}|)$. More details on the relation between MTP and ETN are provided in the original paper by Hodapp and ShapeevHodapp and Shapeev (2024)
This enables learning similarities between the all radial features from the data through the parameter tensors $A$ and $B$, avoiding the problem of an exponentially growing size of the feature vectorsHodapp and Shapeev (2024). As for MTP, the total energy of ETN is the sum of contributions $V^{\rm ETN}({\bf\mathfrak{n}}_{i})$ for $N$ atoms.
The number of coefficients of the ETN potentials is proportional to $d\bar{r}^{2}\bar{n}$, where $\bar{r}$ is the average rank of the tensor network, and $\bar{n}$ is the average size across all dimensions of the feature vector. This makes the growth of their parameter space more controllable than, e.g., for MTPs.
II.3 Fitting
For finding parameters of MLIPs, e.g. MTP or ETN parameters, we fit it on a (quantum-mechanical) training set. Let $K$ be a number of configurations in the training set and $N$ be a number of atoms. Denote a vector of MTP/ETN parameters to be found by ${\bm{\theta}}$. We find the optimal parameters ¯${\bm{\bar{}}{\theta}}$ by solving the following optimization problem (minimization of the objective function)
$\begin{array}[]{c}\displaystyle\sum\limits_{k=1}^{K}\Bigl{[}w_{\rm e}\left(E^{\rm MLIP}_{k}({\bm{\theta}})-E^{\rm QM}_{k}\right)^{2}+\\ \displaystyle w_{\rm f}\sum_{i=1}^{N}\left|{\bf f}^{\rm MLIP}_{i,k}({\bm{\theta}})-{\bf f}^{\rm QM}_{i,k}\right|\Bigr{]}^{2}\to\operatorname{min}.\end{array}$ (9)
We start from randomly initialized MLIP parameters. The optimal parameters ¯${\bm{\bar{}}{\theta}}$ are found numerically, using the iterative method to minimize the non-linear objective function, namely the Broyden-Fletcher-Goldfarb-Shanno algorithm. Thus, after optimization, the parameters ¯${\bm{\bar{}}{\theta}}$ are near the local minimum of the objective function. In this objective function, $E^{\rm QM}_{k}$ and ${\bf f}^{\rm QM}_{i,k}$ are the reference energies and forces, i.e., the ones to which we fit the MLIP energies $E^{\rm MLIP}_{k}$ and forces ${\bf f}^{\rm MLIP}_{i,k}$ and thus optimize the MLIP parameters ${\bm{\theta}}$. The factors $w_{\rm e}$ and $w_{\rm f}$ are non-negative weights which express the importance of energies and forces with respect to each other. We took $w_{\rm e}=1$ and $w_{\rm f}=0.01$. We refer to the minimization problem (9) as MLIP fitting.
II.4 Active Learning
Assume that we found the vector of the optimal MTP parameters ¯${\bm{\bar{}}{\theta}}$ after solving (9) and let the length of the vector be $m$. We then construct a matrix
$\mathsf{B}=\left(\begin{matrix}\frac{\partial E^{\rm MTP}_{1}}{\partial\theta_{1}}({\bm{\bar{\theta}}})&\ldots&\frac{\partial E^{\rm MTP}_{1}}{\partial\theta_{m}}({\bm{\bar{\theta}}})\\ \vdots&&\vdots\\ \frac{\partial E^{\rm MTP}_{K}}{\partial\theta_{1}}({\bm{\bar{\theta}}})&\ldots&\frac{\partial E^{\rm MTP}_{K}}{\partial\theta_{m}}({\bm{\bar{\theta}}})\end{matrix}\right).$
From this matrix, we select a set of the $m$ most linearly independent rows, i.e., we find a submatrix $A$ with the maximum absolute value of the determinant (maximum volume). To find a submatrix $A$ of maximum volume from the matrix $B$, we use the MaxVol algorithmGoreinov et al. (2010) which ensures the diversity of the configurations corresponding to the rows of $A$.
After training the initial MTP and constructing the matrix $A$, we start a molecular dynamics (MD) simulation and compute the extrapolation grade for each configuration ${\bm{x}^{*}}$ created during this simulation
$\gamma({\bm{x}^{*}})=\max\limits_{1\leq j\leq m}|c_{j}|,$ (10)
where the vector $\textbf{c}=(c_{1},\ldots,c_{m})$ for the configuration ${\bm{x}^{*}}$ is computed by
$\textbf{c}=\Bigl{(}\frac{\partial E^{\rm MTP}}{\partial\theta_{1}}(\bm{\bar{\theta}},{\bm{x}}^{*}),\dots,\frac{\partial E^{\rm MTP}}{\partial\theta_{m}}(\bm{\bar{\theta}},{\bm{x}}^{*})\Bigr{)}A^{-1}.$ (11)
We also introduce two thresholds: $\gamma_{\rm save}$ and $\gamma_{\rm break}$. Now we have all the prerequisites to formulate our AL algorithm.
  1. 1.
    Run the MD simulation with extrapolation control. The MD simulation algorithm generates configurations for which energy and forces are computed. An extrapolation grade is also calculated for each configuration. Depending on the degree of extrapolation $\gamma({\bm{x}^{*}})$ there are three possibilities: (i) if $\gamma({\bm{x}^{*}})<\gamma_{\rm save}\approx 2$ then we proceed with the calculation of the energy and forces as we deal with interpolation or insignificant extrapolation; (ii) if $\gamma_{\rm save}<\gamma({\bm{x}^{*}})<\gamma_{\rm break}\approx 10$ then we save the configuration to a file with the extrapolative configurations and proceed with the calculation of the energy and forces; (iii) if $\gamma({\bm{x}^{*}})>\gamma_{\rm break}$ then we break the calculation as the extrapolation degree exceeds the critical threshold $\gamma_{\rm break}$.
  2. 2.
    Selection of configurations. In the previous step, the extrapolative configurations were saved to the file. However, the saved file may contain a large number of similar configurations. According to our active learning strategy, we select among the extrapolative configurations those that maximize the volume of the matrix $A$. Thus, in this step, we select the configurations to be added to the training set.
  3. 3.
    Quantum-mechanical calculations. After selecting the appropriate configurations, we calculate their quantum-mechanical energy and forces.
  4. 4.
    MTP retraining. After ab intio calculations of the configurations, we add them to the training set, retrain MTP, and update the matrix $A$.
We repeat steps 1-4 until no configuration is preselected during MD simulations. We note that the active learning algorithm can only be used with MTP and was implemented only in the MLIP-2 codeNovikov et al. (2020). For the moment, we have not implemented a similar active learning algorithm for ETN.
II.5 Range separation
In the present study, we added the D2Grimme (2006) or D3Grimme et al. (2010) dispersion corrections to the local MTP and ETN potentials. We implemented the D2 correction in both the MLIP-2 and MLIP-4 codes, adding it to both MTP and ETN. The D3 correction was implemented only in the MLIP-4 code and combined only with ETN.
The local potentials, namely MTP and ETN, were trained on configurations with total energies and forces. In contrast, potentials that explicitly incorporate dispersion interactions via D2 or D3, namely, MTP+D2, ETN+D2, ETN+D3, were obtained in the following way. We first trained pure MTP and ETN models on configurations with D2 or D3 contributions subtracted from total energies and forces, and after training, dispersion correction was explicitly added to the top of a trained MLIP.
The explicit incorporation of dispersion via D2 or D3 allows to maintain a cutoff radius of 5—6 Å for the machine learning part of these potentials, while the cutoff radius for D2 or D3 is set to 15 Å for all simulations presented in this work.
II.6 D2 correction
The D2 dispersion correction was introduced by Grimme et al. Grimme (2006) to account for dispersion interactions that are not inherently included in the DFT framework.
D2 considers only the two-body dipole-dipole contribution to the dispersion energy, leading to the following form of energy expression:
$E^{D2}=-s_{6}\sum_{ij}\frac{C_{6}^{ij}}{r_{ij}^{6}}f_{\rm damp}(r_{ij}),$ (12)
where $i,j$ are indices of atoms, $s_{6}$ is a scaling factor dependent on a particular density functional, $C_{6}^{ij}$ is a coefficient describing interactions between atoms $i$ and $j$, $r_{ij}$ is an interatomic distance between atoms $i$ and $j$, and $f_{\rm damp}(r_{ij})$ is a damping function.
In this scheme, the coefficients $C_{6}^{ii}$ are precomputed parameters and are calculated from the empirical formula:
$C_{6}^{ii}=0.05NI_{A}\alpha_{A}^{0},$ (13)
where $N$ is a scaling factor, $I_{A}$ is the first ionization potential calculated with PBE0, and $\alpha_{A}^{0}$ is static dipole polarizability. The $C_{6}^{ij}$ are defined as the geometric mean of $C_{6}^{ii}$ and $C_{6}^{jj}$:
$C_{6}^{ij}=\sqrt{C_{6}^{ii}C_{6}^{jj}}.$ (14)
The Fermi damping function $f^{\rm Fermi}_{\rm damp}$ is utilized in D2:
$f^{\rm Fermi}_{\rm damp}=\frac{1}{1+e^{{20(r_{ij}/(s_{R}R_{\rm vdW})-1)}}},$ (15)
where $s_{R}$ is a scaling parameter, $R_{\rm vdW}$ is a sum of van der Waals radii of interacting atoms.
However, in our implementation of the D2 correction (as well as the D3 correction), we chose Becke-Johnson damping function $f^{(n)}_{\rm damp,BJ}$:
$f^{(n)}_{\rm damp,BJ}=\frac{r^{n}}{r^{n}+(a_{1}R_{0}+a_{2})^{n}}$ (16)
with $n=6$. We adjusted the coefficients $a_{1}$, $a_{2}$, and $R_{0}$ on CCl4, CH4, and toluene dimer binding curves.
II.7 D3 correction
In contrast to the D2 scheme, the D3 dispersion correction Grimme et al. (2010) explicitly takes into account information about the atomic environment via fractional coordination number. Moreover, in the D3 approach, the dispersion energy includes not only the dipole-dipole interaction term, proportional to $1/r^{6}$, but also the higher order dipole-quadrupole contribution proportional to $1/r^{8}$:
$E^{D3}=-\sum_{ij}\sum_{n=6,8}s_{n}\frac{C_{n}^{ij}}{r_{ij}^{n}}f_{\rm damp}^{(n)}(r_{ij}),$ (17)
where $i,j$ are the indices of atoms, $s_{n}$ is a scaling factor dependent on the density functional, $C_{n}^{ij}$ is a coefficient describing interactions between atoms $i$ and $j$, $r_{ij}$ is the interatomic distance between atoms $i$ and $j$, and $f_{\rm damp}^{(n)}(r_{ij})$ is damping function, dependent on the power $n$.
In the D3 correction, $C_{n}^{ij}$ explicitly includes information on the environments of both $i$ and $j$ atoms via coordination numbers:
$CN^{i}=\sum_{i\neq j}^{N_{\rm atoms}}\frac{1}{1+e^{{-16(4(R_{i,{\rm cov}}+R_{j,{\rm cov}})/(3r_{ij})-1)}}},$ (18)
where $R_{i,{\rm cov}}$, $R_{j,{\rm cov}}$ are covalent radii of elementsPyykkö and Atsumi (2009).
The computed coordination numbers are further used for interpolating $C_{6}^{ij}$ based on a set of precomputed coefficients $C_{6,\rm ref}^{ij}$, obtained for reference compoundsGrimme et al. (2010):
$C_{6}^{ij}(CN^{i},CN^{j})=\frac{\sum_{p}^{N_{i}}\sum_{q}^{N_{j}}C_{6,\rm ref}^{ij}(CN_{p}^{i},CN_{q}^{j})L_{pq}}{\sum_{p}^{N_{i}}\sum_{q}^{N_{j}}L_{pq}},$ (19)
where $L_{pq}$ is defined as:
$L_{pq}=e^{{-4\big{(}(CN^{i}-CN^{i}_{p})^{2}+(CN^{j}-CN^{j}_{q})^{2}\big{)}}}.$ (20)
Based on the obtained $C_{6}^{ij}$, the coefficients $C_{8}^{ij}$ for the higher-order term in dispersion energy (17) is computed:
$C_{8}^{ij}=3C_{6}^{ij}\sqrt{Q_{i}Q_{j}},$ (21)
where $Q_{i}$, $Q_{j}$ are precomputed parameters derived from atomic densities and atomic numbers of elements $i$ and $j$Grimme et al. (2010).
For the D3 correction, two types of damping function can be utilized: zero-damping and Becke-Johnson damping (BJ) (16). For our implementation of D3 we chose the BJ damping with $n=6$ for the $1/r^{6}$ damping term, and $n=8$ for the higher order term.
II.8 Protocol of datasets generation and training of MLIPs
We collected training sets with total energies and forces via the AL approach described above and implemented in the MLIP-2 code for MTP. For fitting of models that explicitly incorporate dispersion via D2 or D3 (MTP+D2, ETN+D2, ETN+D3), the training set with total energies and forces was post-processed by subtracting D2 or D3 contributions from the total energies and forces. The general workflow for obtaining training sets for each of the considered models is shown in Figure 1.
Figure 1: The general scheme for generating training sets for all the MLIPs considered. At the first step, the AIMD simulation is conducted and initial MTP is trained on the configurations uniformly chosen from the AIMD trajectory. Next, a pretraining set formed from uniformly selected configurations and configurations, additionally selected from the AIMD trajectories with the MaxVol algorithmPodryabinkin and Shapeev (2016). The initial MTP retrained on the resulting pretraining set. After that the AL of MTP during MD simulations is carried out, and results in the actively selected training set with total energies and forces. Further steps correspond to training of MTP and ETN on the dataset including total energies and forces, and subtracting D2 or D3 contributions and utilizing the resulting training sets for fitting of MTP or ETN part of MTP+D2, ETN+D2, ETN+D3 potentials.
We started creating a training set from ab initio molecular dynamics (AIMD) simulations. An initial configuration was generated using Packmol Martínez et al. (2009). Next, we uniformly chose about 20 configurations from this AIMD trajectory to avoid correlations between configurations and fit an initial MTP. After that, we selected configurations using the MaxVol algorithm from the pool of configurations generated during AIMD, updated the initial training set, and refitted the MTP, obtaining pretrained MTP. Subsequently, we actively trained the pretrained MTP during 10 parallel MD simulations of 100 ps each using the AL procedure described above, utilizing LAMMPS softwareThompson et al. (2022) for MD. As a result, we obtained the actively selected training set with total energies and forces. During both AIMD and MD in the AL procedure, a timestep was set to 1 fs.
The collected training set was then used to train the MTP and ETN models. After subtracting the contributions of D2 or D3 from the total energies and forces in the collected training set, we obtained training sets for the MTP+D2, ETN+D2, and ETN+D3 models.
For all the investigated systems, the level of the MTP-based potentials (MTP and MTP+D2) was set to 16, corresponding to 222 parameters to be optimized during the fitting, and for the ETN-based models (ETN, ETN+D2, and ETN+D3) we took 124 coefficients to be fitted. We note that the AL procedure was performed only for MTP and the resulting training set may not be optimal for other tested MLIPs, especially for the ETN-based models. We emphasize that in this procedure we implicitly assume, since ETN has fewer parameters, that the training set actively selected by MTP would also be suitable for ETN.
For creating validation sets, the configurations were uniformly sampled from independent 100 ps MD simulation conducted with actively trained MTP. The energies and forces for the sampled configurations were then computed using ab initio calculations. The ratio between the size of the training and validation sets was set to 8:2.
II.9 Assessing the quality of MLIPs
To compare the quality of the fitted MLIPs we evaluated training and validation errors of predicting energies and forces. In addition, we designed two series of tests. In the first series, we evaluated the accuracy of potentials in predicting intermolecular interactions. To that end, we predicted binding curves for a diverse set of dimers and compared them with the ones obtained with ab initio calculations. We calculated the deviation between the position and depth of the minima, as well as the full width of the well at half its depth (referred to as full-width at half-minimum or FWHM) obtained with potentials and ab initio calculations (see Figure 2). In this test, the binding curves were represented as dependencies of energy on the distance between the centers of mass of individual molecules. The second series of tests aimed to assess the accuracy of density and RDF calculated during MD, performed with LAMMPS softwareThompson et al. (2022). We used experimental data as reference.
Figure 2: Error metrics proposed for assessing accuracy of MLIPs in predicting dimer binding curves.
To ensure that the error estimations are statistically meaningful, we conducted each test for an ensemble of five potentials for each type of MLIP model considered, except for calculating density and RDF with ETN+D3, as this type of MLIP is more computationally demanding than others.
III Results and discussion
In this section, we present results of application of the described MLIPs to the modeling of liquid carbon tetrachloride, liquid methane, and liquid toluene. For all systems, we focused on evaluating the accuracy differences between MLIP(l)’s and MLIP+D’s.
We note that for carbon tetrachloride, we performed two test cases. In the first case, we compared the accuracy of MLIP(l)’s and MLIP+D’s when the $R_{\rm cut}$ for all potentials is set to 5.5 Å, which limits the ability of MLIP(l)’s to fully capture the range of dispersion interactions. In the second case, we explore to what extent the accuracy of MLIP(l)’s can be improved by increasing the $R_{\rm cut}$ to 7.5 Å, allowing them to fully capture the range of dispersion interactions (while maintaining a cutoff radius of 5—6 Å for MLIP+D’s). For methane and toluene, we took only the $R_{\rm cut}$ of 7.5 Å for the MLIP(l)’s and of 6 Å for MLIP+D’s.
III.1 CCl4 modeling with $R_{\rm cut}=5.5\text{\AA}$
III.1.1 Datasets generation
To create the pretraining set, we performed AIMD at 298 K for 1 ps in the NVT ensemble. In the subsequent AL process, new configurations were sampled during MD conducted in the NpT ensemble at 293 K and 1 bar. The accumulated training set consisted of 478 configurations, each consisting of 6 CCl4 molecules with a simulation cell size ranging from 10 to 12 Å.
Ab initio simulations were performed using the CP2K package Kühne et al. (2020). We used the Gaussian Plane Wave (GPW) approach Lippert, Hutter, and Parrinello (1997) with QZV2P-GTH-q4 and QZV2P-GTH-q7 basis sets, along with GTH-PBE-q4 and GTH-PBE-q7 pseudopotentials for carbon and chlorine atoms, respectively. The GPW calculation also requires the convergence of the real-space integration grid by adjusting the energy cutoff VandeVondele et al. (2005), which was set to 400 Ry. All calculations were performed at the PBE-D3 level of theory.
III.1.2 Training and validation
For each of the MLIP models, an ensemble of five potentials was trained. The training and validation root mean square errors (RMSEs), averaged over the ensembles, are presented in Figure 3. They indicate an improvement in energy accuracy when moving from MTP to MTP+D2, while the difference between ETN and ETN+D2, ETN+D3 is not clearly distinguishable.
Figure 3: Training and validation RMSEs with one-$\sigma$ confidence interval on liquid CCl4 configurations obtained with MLIPs with $R_{\rm cut}=5.5\text{\AA}$.
III.1.3 Binding curves
In this subsection, we predict dimer binding curves for investigating possible sources of training and validation errors, originating from intermolecular interactions. For this test, we considered a diverse set of dimers (Figure 4) to evaluate the precision across a range of interaction strengths and distances (Table 1).
Figure 4: Dimers of CCl4 used for generation of the binding curves. The labels correspond to description of mutual orientation, based on treatment of CCl4 molecule as tetrahedron. F corresponds to “face”, V – “vertex”, E – “edge”, and, finally, s and e correspond to staggered and eclipsed conformations.
FFs FFe FVs FVe EEs EEe EF EV VVs VVe
Emin, meV -95 -74 -66 -66 -44 -72 -73 -62 -21 -18
rmin, Å 4.9 5.2 5.9 5.9 6.0 5.5 5.4 6.2 7.3 7.3
Table 1: Depths and positions of binding curves minima for carbon tetrachloride dimers, calculated with PBE-D3 using CP2K.
The mean absolute percentage error (MAPE) and the RMSE averaged over the dimers and ensembles of MLIPs are illustrated in Figure 5(a). The accuracy in the energy at the minimum improves from ETN to ETN+D2 and further to ETN+D3, while for MTP and MTP+D2 the difference is minor. In addition, MLIP+D’s appear to have a slightly higher accuracy in predicting FWHM than MLIP(l)’s.
Figure 5: Errors of predicting the minima of binding curves: depth (E), position (r) and full-width at half-minimum (FWHM). a) MAPE and RMSE, averaged over ensembles of potentials and dimers; b) MAPE for each of the dimers considered, averaged over ensembles of potentials.
MAPE also indicates that predicting the depth and FWHM is more challenging than predicting the position of the minimum. Thus, MLIP+D’s predict FWHM and minima depth with approximately 20-40% error. Furthermore, the precision is not uniform across different dimers (Figure 5(b)) with the highest error corresponding to the VVs and VVe dimers having the weakest interaction and the farthest minimum position (Table 1). According to these tests, MTP shows slightly better accuracy than ETN, while ETN+D2 and ETN+D3 are more precise than MTP+D2. Additionally, ETN+D3 slightly outperforms other MLIP+D’s.
The mean bias error (MBE) (Figure 6(a)) averaged over the ensembles of MLIPs and dimers demonstrates that all the potentials excluding ETN+D3 tend to underestimate the FWHM, especially the MTP and ETN potentials. However, in contrast to MTP, the ETN potential additionally tends to significantly overbind the dimers. This can be demonstrated on the binding curves, averaged over the ensembles of MLIPs (Figure 6(b)) and presented together with a one-$\sigma$ confidence interval.
Figure 6: a) Mean bias error (MBE) on binding curves averaged over ensembles of models and dimers; b) binding curves for the FFs, EEe, VVe dimers, calculated with PBE-D3 and predicted with MLIPs. The averaged curves (dashed line) are plotted together with the one-$\sigma$ confidence interval.
The MLIP(l)’s not only exhibit a larger uncertainty but also do not have the correct $1/r^{6}$ asymptotic behavior. The asymptotic behavior slightly improves when dispersion interactions are included via D2, however there is an observable difference compared to the ab initio binding curve, likely due to the difference between D2 $C_{6}$ coefficients used in the dispersion correction of MLIPs and D3 $C_{6}$ coefficients used in ab initio calculations. The best asymptotic behavior is achieved by ETN+D3.
III.1.4 Molecular Dynamics
We proceeded with MD simulations to quantify the impact of differences between MLIPs observed in the binding curves test on density and RDF. For density estimation, we performed 500 ps MD in the NpT ensemble at 293 K and 1 atm with a timestep of 1 fs. The initial simulation cell had a size of 30 Å and consisted of 162 molecules. The resulting density value was averaged over the last 100 ps of the simulation. The dependency of density on time is shown in Figure 7(a), and the MAPE compared to the experimentHien, Ponter, and Peier (1978) is presented in Table 2.
Figure 7: Predicted with the ensembles of MLIPs: a) dependencies of density on time together with one-$\sigma$ confidence intervals; b) average RDFs together with two-$\sigma$ confidence intervals (the larger confidence interval was chosen for illustrative purposes).
In Figure 7(a) we see that the MTP potential does not maintain stable molecular dynamics in the liquid phase, while this issue is resolved by explicit incorporation of dispersion interactions in MTP+D2, resulting in a 25.6% deviation from the experimentally measured value of 1.594 g/cm3. Unlike MTP, the ETN potential is capable of preserving the liquid phase, likely due to the overbinding of molecules, as demonstrated in the dimer binding curves test. Meanwhile, ETN+D2 and ETN+D3 appear to have higher accuracy compared to ETN, achieving a deviation of approximately 22% from the experimental density.
MTP MTP+D2 ETN ETN+D2 ETN+D3
MAPE 0.991 0.256 0.256 0.216 0.222
$\pm$ 0.001 $\pm$ 0.035 $\pm$ 0.044 $\pm$ 0.009 $\pm$ 0.015
Table 2: Density MAPE and its one-$\sigma$ confidence interval, compared to experimentHien, Ponter, and Peier (1978).
To calculate the RDF, we performed MD starting from the simulation cell with the previously converged density, following a 10 ps temperature equilibration in the NVT ensemble at 293 K. The RDF was averaged over 100 ps in the NVE ensemble. The intensities and positions of the maxima of the calculated RDFs for carbon atoms were compared with those obtained experimentallyLowden and Chandler (1974).
In Figure 7(b) the predicted RDF, averaged over the ensemble of MLIPs, is illustrated together with a two-$\sigma$ confidence interval. All MLIPs produce nearly equivalent RDFs (Table 3), at the same time exhibiting a shift compared to the experimental data, likely due to an underestimation of the density.
MTP+D2 ETN ETN+D2 ETN+D3
MAPE(g) 0.041 0.065 0.056 0.056
$\pm$ 0.034 $\pm$ 0.040 $\pm$ 0.053 $\pm$ 0.053
MAPE(r) 0.036 0.031 0.033 0.033
$\pm$ 0.017 $\pm$ 0.019 $\pm$ 0.020 $\pm$ 0.020
Table 3: MAPE and two-$\sigma$ confidence interval of positions and intensities of RDF maxima, compared to experimentLowden and Chandler (1974).
In contrast to the accuracy obtained from the binding curve test, ETN+D3 did not outperform ETN+D2 in predicting density and RDF.
III.2 CCl4 modeling with extended $R_{\rm cut}$
In this subsection, we test MLIP(l)’s with the $R_{\rm cut}$ extended to 7.5 Å, while the $R_{\rm cut}$ of MLIP+D’s is set to 6 Å. In this manner, we aim to investigate the ability of MLIP(l)’s to describe both covalent and dispersion interactions simultaneously, which may be challenging because the dispersion interactions are orders of magnitude weaker than intramolecular interactions and are comparable to the magnitude of MLIP fitting errors.
III.2.1 Datasets generation
For ab initio calculations, we used the VASP softwareKresse and Furthmüller (1996), and chose PAW_PBE pseudopotentials, a 900 eV kinetic energy cutoff, and a k-point grid including only gamma-point. All calculations were performed at the PBE-D3 level of theory.
The procedure for generating the training set with total energies and forces was similar to that described earlier. The AIMD of 500 ps was carried out in the NVE ensemble with positions generated by PackmolMartínez et al. (2009) and velocities initialized at 298 K. The MD conditions in the AL process were the same as those used before. At the end of the AL procedure, the accumulated training set comprised 805 configurations, each containing 11 molecules with a cell size ranging from 15 to 17 Å.
III.2.2 Training and validation
Following the previously described procedure, we trained an ensemble of five potentials for each MLIP model. Training and validation errors are presented in Figure 8 and demonstrate the first impact of increasing $R_{\rm cut}$. From the figure we see that the errors of MLIP(l) decreased (by approximately a factor of three for energy and 20% for forces) and reached the level of MLIP+D errors.
Figure 8: Training and validation RMSEs with one-$\sigma$ confidence interval on liquid CCl4 configurations obtained with MLIP(l)’s with $R_{\rm cut}=7.5\text{\AA}$ and MLIP+D’s with $R_{\rm cut}=6\text{\AA}$.
III.2.3 Binding curves
For the present test, we used the same set of dimers as was introduced above (Figure 4). We note that the binding curves were recalculated with VASP. The results are presented in Table 4, demonstrating minor changes in interaction energies and minima positions.
FFs FFe FVs FVe EEs EEe EF EV VVs VVe
Emin, meV -93 -75 -63 -62 -48 -69 -71 -59 -20 -21
rmin 4.9 5.2 5.9 5.9 6.0 5.5 5.4 6.2 7.2 7.2
Table 4: Depths and positions of binding curves minima for carbon tetrachloride dimers, calculated with PBE-D3 using VASP.
Similarly to the training and validation errors, energy MAPE and RMSE, obtained for the prediction of the binding curves (Figure 9(a)), decreased by approximately 20 and 15%, respectively, compared to the errors obtained for MLIPs with a smaller $R_{\text{cut}}$ However, no significant differences in accuracy were observed between MLIP(l)’s and MLIP+D’s, except for the reduction of errors and the uncertainty of the FWHM estimation.
Figure 9: Errors of describing the minima of binding curves: depth (E), position (r) and full-width at half-minimum (FWHM). a) MAPE and RMSE, averaged over ensembles of potentials and dimers; b) MAPE for each of the dimers considered, averaged over ensembles of potentials.
From Figure 9(b) we see that the errors still depend on the mutual orientation of the molecules. Specifically, the energy MAPE for VVs and VVe dimers remains significantly higher than that for others. In addition, MLIP(l)’s now demonstrate smaller energy errors, while possessing higher errors in predicting FWHM on these dimers compared to MLIP+D’s.
Unlike MLIPs with smaller $R_{\rm cut}$, in the present case we did not observe a systematic hardening of the binding curves, as indicated by the MBE (Figure 10(a)). Specifically, only MTP tends to soften the binding curves, while other MLIPs demonstrate a low MBE together with lower MAPE and RMSE (Figure 9 (a)), indicating an accurate prediction of FWHM. The energy error has also been reduced, and most of the MLIPs tend to slightly underestimate the depth of the minima, while ETN continues to show a preference for overbinding.
Figure 10: a) MBE on binding curves averaged over ensembles of models and dimers; b) binding curves for the FFs, EEe, VVe dimers, calculated with PBE-D3 and predicted with MLIPs. The averaged curves (dashed line) are plotted together with the one-$\sigma$ confidence interval.
The large errors in predicting FWHM with the MTP potential might result from less accurate asymptotics compared to other MLIPs. This leads to the existence of unphysical minima on the VVe dimer binding curve (Figure 10(b)).
We note that in the present case, ETN+D3 did not exhibit a higher accuracy than ETN+D2, while being more computationally expensive. Moreover, it did not provide more accurate density and RDF estimations when MLIPs with smaller $R_{\rm cut}$ were tested. For these reasons, we do not perform MD simulations for ETN+D3 in the present case.
III.2.4 Molecular Dynamics
The simulation conditions were the same as for the previous MD tests, while the initial configuration size was 30 Å and it consisted of 168 molecules.
Tables 5 and 6 show the errors in density and RDF calculated by different potentials. Similarly to the binding curves, the errors of all MLIPs are close to each other.
MTP MTP+D2 ETN ETN+D2
MAPE 0.151 0.204 0.170 0.177
$\pm$ 0.011 $\pm$ 0.016 $\pm$0.008 $\pm$ 0.010
Table 5: Density MAPE and its one-$\sigma$ confidence interval, compared to experimentHien, Ponter, and Peier (1978).
MTP MTP+D2 ETN ETN+D2
MAPE(g) 0.099 0.062 0.118 0.126
$\pm$ 0.075 $\pm$ 0.051 $\pm$ 0.088 $\pm$ 0.094
MAPE(r) 0.028 0.030 0.027 0.029
$\pm$ 0.013 $\pm$ 0.017 $\pm$ 0.015 $\pm$ 0.013
Table 6: MAPE and one-$\sigma$ confidence interval of positions and intensities of RDF maxima, compared to experimentLowden and Chandler (1974).
Thus, we demonstrated that it is sufficient to introduce an $R_{\rm cut}$ capturing the range of dispersion interactions for MLIP(l)’s for accurate modeling of liquid carbon tetrachloride, since the difference in accuracy between the MLIP(l) and MLIP+D potentials is minor for all properties calculated excluding FWHM.
However, we want to emphasize that the MLIP+D models still have advantages for the extended-cutoff MLIP(l)’s. Indeed, employing a smaller $R_{\rm cut}$ while incorporating explicit dispersion corrections enhances the computational efficiency of potential fitting and subsequent simulations. This efficiency stems from the reduced computational cost of DFT calculations for smaller configurations, as well as the faster training and evaluation of energies and forces with potentials that have a smaller cutoff radius.
In view of all the results for CCl4, for the other two systems we decided to limit the tests for MLIP+D’s with $R_{\rm cut}=6\text{\AA}$ and MLIP(l)’s with $R_{\rm cut}=7.5\text{\AA}$.
III.3 CH4 modeling
The molecular structure of methane is essentially the same as that of carbon tetrachloride, suggesting the ability of MLIP(l)’s to accurately learn dispersion interactions. However, in liquid methane the dispersion interactions are weaker than in carbon tetrachloride. This raises the question of whether MLIP(l)’s can accurately capture dispersion interactions together with intramolecular interactions when the difference in their magnitudes increases.
III.3.1 Datasets generation
For liquid methane modeling, all ab initio calculations were performed using VASP, with the same settings as those used for CCl4.
AIMD simulations to collect the pretraining set were performed in the NVE ensemble for 500 fs starting from the configuration generated with PackmolMartínez et al. (2009), with velocities initialized at 298 K. During the AL process, MD simulations were performed at 150 K and 10 bar. At the end of the AL process, the dataset accumulated 306 configurations, each comprising 57 molecules with simulation box sizes ranging from 15 to 18 Å.
III.3.2 Training and validation
As in previous cases, for each MLIP model, an ensemble of five potentials was trained, and the RMSEs averaged over the ensembles are presented in Figure 11. The magnitude of the errors is about four times lower for energies and two times lower for forces than that obtained for CCl4 with MLIPs with extended $R_{\rm cut}$. Similarly to CCl4, no significant differences are observed between MLIP(l)’s and MLIP+D’s, excluding thecg improvement of accuracy from ETN to ETN+D2 and further to ETN+D3.
Figure 11: Training and validation RMSEs with one-$\sigma$ confidence interval on liquid CH4 configurations.
III.3.3 Binding curves
We constructed binding curves for the same types of dimers as for CCl4 (Figure 4). The ab initio energies and positions of the minima are presented in Table 7. Compared to carbon tetrachloride, CH4 dimers exhibit closer and shallower minima, furthermore, the interaction energies are within a narrower range compared to CCl4 dimers.
FFs FFe FVs FVe EEs EEe EF EV VVs VVe
Emin, meV -35 -33 -29 -29 -25 -29 -31 -25 -17 -18
rmin 3.7 3.8 4.1 4.1 4.0 4.0 3.9 4.2 4.6 4.6
Table 7: Depths and positions of binding curves minima for methane dimers.
Figure 12(a) shows the binding curve errors, averaged over the ensembles and dimers. There we observe that the precision in energy of MLIP(l)’s and MLIP+D’s is nearly equal, while the ETN family exhibits higher accuracy than the MTP family. However, similar to the prediction of the CCl4 binding curves (Figure 9), MLIP+D’s still have an advantage having considerably lower RMSE and uncertainty in predicting the FWHM.
Figure 12: Errors of describing the minima of binding curves: depth (E), position (r) and full-width at half-minimum (FWHM). a) MAPE and RMSE, averaged over ensembles of potentials and dimers; b) MAPE for each of the dimers considered, averaged over ensembles of potentials.
In contrast to CCl4 binding curves, CH4 errors are not strongly dependent on the mutual orientation of the molecules, showing a uniformly high level of accuracy. This may be attributed to the narrower range of interaction energies in CH4 dimers, which does not result in the discrimination of certain dimers, as seen in the case of CCl4.
According to MBE, all MLIPs tend to underestimate the depth of the minima (Figure 13(a)), while both MTP and ETN also show a significant overestimation of the potential width. However, no unphysical minima are observed for MTP, as in CCl4 binding curves (Figure 13(b)). This suggests that all MLIPs are qualitatively correct and reached a close level of accuracy in this test.
Figure 13: a) MBE on binding curves averaged over ensembles of models and dimers; b) binding curves for the FFs, EEe, VVe dimers, calculated with PBE-D3 and predicted with MLIPs. The averaged curves (dashed line) are plotted together with the one-$\sigma$ confidence interval.
Similarly to the CCl4 results with the extended cutoff radius of MLIPs, the ETN+D3 does not outperform ETN+D2 and thus will not be used for further simulations due to being more computationally expensive.
III.3.4 Molecular Dynamics
To predict density, a 30 Å cubic cell, consisting of 456 molecules, was equilibrated for 300 ps in the NpT ensemble, at $T=110$ K and $p=20.75$ bar. The resulting density value was averaged over the last 100 ps and compared to the experimental valueHarris and Trappeniers (1980) of 0.4256 g/cm3. The calculated MAPE is presented in Table 8. The errors obtained for all MLIPs are small and close in magnitude.
MTP MTP+D2 ETN ETN+D2
MAPE 0.033 0.045 0.035 0.043
$\pm$ 0.004 $\pm$ 0.005 $\pm$ 0.003 $\pm$ 0.002
Table 8: Density MAPE and its one-$\sigma$ confidence interval, compared to experimentHarris and Trappeniers (1980).
To predict RDFs, the NpT simulation was performed for 150 ps at 92 K and 0.13 bar; subsequently, the ensemble was switched to NVT and the temperature was equilibrated for 10 ps, then the RDF was averaged over 100 ps in the NVE simulation. The resulting deviations of the RDFs from the experimentally observedOobatake, Hayashi, and Machida (1990) are reported in Table 9 and are nearly the same for all MLIPs, similar to the results obtained for CCl4.
MTP MTP+D2 ETN ETN+D2
MAPE(g) 0.036 0.027 0.035 0.031
$\pm$ 0.017 $\pm$ 0.012 $\pm$ 0.010 $\pm$ 0.011
MAPE(r) 0.019 0.019 0.019 0.025
$\pm$ 0.018 $\pm$ 0.011 $\pm$ 0.016 $\pm$ 0.015
Table 9: MAPE and one-$\sigma$ confidence interval of positions and intensities of RDF maxima, compared to experimentOobatake, Hayashi, and Machida (1990).
Note that the pressure conditions of the simulations for density and RDF calculations differ significantly from those during the MD simulations in the AL process, yet the results are highly accurate, which indicates excellent transferability of the trained MLIPs.
III.4 Toluene modeling
As was demonstrated for liquid methane and carbon tetrachloride, MLIP(l) with extended $R_{\rm cut}$ can accurately capture dispersion interactions. However, methane and carbon tetrachloride molecules possess high symmetry, and the covalent and dispersion interactions are well separated in the interatomic distance scale. Now we proceed with the testing of MLIPs on liquid toluene, where the molecule is less symmetrical, the dispersion interactions are strongly dependent on the mutual orientation of molecules, are not clearly separated in the interatomic distance scale.
III.4.1 Datasets generation
All ab initio simulations for toluene were performed using FHI-aimsBlum et al. (2009) software with intermediate settings for basis set and integration grids for both atomic types and a 2$\times$2$\times$2 k-point grid. All simulations were carried out at the PBE-TS level of theory.
The AIMD simulation was performed in the NVT ensemble for 400 fs, at 298 K. The MD conditions in the AL were set to 298 K and 1 atm. At the end of the AL, the training set accumulated 520 configurations with sizes ranging from 15 to 18 Å, each consisting of 13 molecules.
III.4.2 Training and validation
Training and validation errors averaged over ensembles of MLIPs are shown in Figure 14. Judging by the energy validation errors, all the MLIP+D models are more accurate than their local counterparts, while this trend is not observed when looking at force errors.
Figure 14: Training and validation errors obtained for liquid toluene configurations.
III.4.3 Binding curves
The dimers proposed for this test are illustrated in Figure 15. They include dimers with and without stacking interactions, t-shaped dimers, and dimers with molecules lying in the same plane.
Figure 15: Dimers of toluene used for constructing binding curves and minima depths and positions. The labels denote mutual orientations of molecules: "P" – parallel, "T" – t-shaped, "IP" – in plane.
MAPE and RMSE, averaged over the ensembles of MLIP and dimers, are presented in Figure 16(a). The errors obtained are significantly higher than for other systems studied, including the modeling of CCl4 with MLIPs with small $R_{\rm{cut}}$. At the same time, there is a considerable improvement in minimum energy and FWHM accuracy for MLIP+D’s compared to MLIP(l)’s, especially for the ETN family of potentials.
Figure 16: Errors of describing the minima of binding curves: depth (E), position (r) and full-width at half-minimum (FWHM). a) MAPE and RMSE, averaged over ensembles of potentials and dimers; b) MAPE for each of the dimers considered, averaged over ensembles of potentials.
The errors are strongly dependent on the mutual orientation of the molecules (Figure 16(b)), even to a greater extent than for CCl4. The highest MAPEs correspond to the most strongly interacting dimers with staking interactions: P-1, P-2, and P-3. In addition, IP-1, IP-2, and IP-3 also exhibit errors higher than average.
Figure 17: MBE for predicting binding curves minima depth (E), position (r) and full-width at half-minimum (FWHM), averaged over ensembles of potentials and dimers.
According to MBE (Figure 17), the ETN family of potentials tends to strongly overbind dimers, while MTP-based potentials show moderate underbinding. At the same time, the ETN-based potentials significantly underestimate the FWHM, while the MTP-based ones give more accurate predictions. Similarly to CCl4 and CH4, both MLIP(l)’s and MLIP+D’s accurately predict the position of the minima.
The large errors observed for the ETN family of potentials result from strong unphysical oscillations (Figure 18). This behavior is observed for all the MLIPs, but is more severe in the case of the ETN-based ones. Those oscillations are weaker for MLIP+D’s compared to MLIP(l)’s.
Figure 18: Binding curves of P-2, T-1 and IP-1 toluene dimers, predicted with PBE-TS (solid blue line) and with MLIPs (dashed line); the predicted binding curves are averaged over ensembles of MLIPs, and are demonstrated together with one-$\sigma$ confidence interval. All binding curves are plotted as functions of the distance between centers of mass (CM) of individual molecules.
Similarly to CCl4 and CH4, ETN+D3 does not outperform ETN+D2 in this test and, for this reason, will not be used in further simulations.
III.4.4 Molecular Dynamics
For MD simulations, a 30 Å simulation box consisting of 104 molecules was used. The density was equilibrated in the NpT ensemble for 200 ps with a 1 fs time step, at 298 K and 1 atm. The resulting density value was obtained by averaging over the last 100 ps and compared to the experimental valueHarris (2000) of 0.862 g/cm3. Figure 19(a) presents the density MAPE which is significantly smaller for MLIP+D’s than for MLIP(l)’s.
The density as a function of time is illustrated in Figure 19(b) where the trajectories averaged over ensembles of MLIPs are shown together with the one-$\sigma$ confidence intervals. The density predicted by ETN exhibits a large uncertainty, indicating large deviations in the MLIP training results. Furthermore, the mean density evaluated with ETN significantly overestimates the experimental value. This is likely a consequence of the strong overbinding of molecules observed in the binding curve test (Figure 17). For the ETN + D2 potential, these issues are resolved, and a density value closer to the experiment is achieved together with a lower spread. MTP also shows higher error and uncertainty compared to MTP+D2, although the difference is not as dramatic as for ETN and ETN+D2.
Figure 19: a) Density MAPE and its one-$\sigma$ confidence interval, compared to experiment; b) density as a function of time together with one-$\sigma$ confidence interval; c) predicted RDFs (green and brown lines) and experimental RDFs (solid blue line); obtained for C7 (methyl) and C4 (opposite to the methyl group) carbons, averaged over ensemble of MLIPs, together with one-$\sigma$ confidence intervals.
For RDF calculation, a 10 ps equilibration in the NVT ensemble was conducted for a simulation cell with the converged density from the previous simulation. After this, the ensemble was switched to NVE and 100 ps MD was performed with a 1 fs time step. During the NVE run, the configurations were collected every 50 steps and post-processed using MDAnalysisMichaud-Agrawal et al. (2011). The RDFs were calculated based on the C4 carbon atoms (opposite to the methyl group) and the C7 carbon atoms (methyl carbon). Three types of RDF were computed: C7–C4, C4–C4, and C7–C7. Note that in the C7-C4 RDF the intramolecular contribution was subtracted. The resulting RDFs were compared with the experimentFalkowska et al. (2016) and presented in Figure 19(c).
RDFs obtained with all the MLIPs, except ETN, show qualitative agreement with the experiment. In contrast, the average C4–C4 RDF predicted with ETN has a false maximum. In addition, individual RDFs predicted by individual ETN potentials in the ensemble demonstrate significant spread from each other, leading to broad uncertainty regions. For the ETN+D2 potential, the width of the uncertainty region is substantially reduced and a better agreement with the experiment is achieved. The difference between MTP and MTP+D2 is less pronounced than for ETN and ETN+D2. However, a reduction in the uncertainty region is still observed, along with an improved alignment between the mean and experimental RDFs.
IV Conclusions
In this work, we considered the problem of using MLIPs to model systems with strong long-range dispersion interactions. A typical MLIP cutoff radius is 5–6 Å which is much less than the range of dispersion interactions, hence the idea to include explicit dispersion interactions on top of a general-purpose MLIP.
We studied the impact of explicit incorporation of dispersion interactions in MTPs and ETN potentials via the D2 and D3 correction schemes. The impact was measured based on training and validation errors, errors in predicting binding curves, density, and RDFs of a liquid phase of three systems—carbon tetrachloride (CCl4), methane (CH4), and toluene.
Our first finding is that the accuracy of a typical MLIP (with a cutoff radius of 5.5 Å) is significantly improved after incorporating explicit dispersion. We observe an improvement in accuracy for predicting energy/forces and binding curves, and this improvement consistently translates into a higher accuracy on the finite-temperature properties.
Comparing the D2 and D3 corrections, we found that the incorporation of D2 is sufficient, as it provides the same order of accuracy as D3 in the systems we tested. This finding suggests that D2 can be a practical choice for many applications, as D3 is more computationally expensive (or at least harder to implement efficiently).
Another finding is that for the two relatively simple systems, CCl4 and CH4, the improvement in accuracy can instead be achieved by simply increasing the cutoff radius of the local MTP and ETN to $7.5\text{\AA}$, except for the full-width at the half-minimum metrics, in which the MLIP+D’s are still more accurate. We note that we did not systematically study the effect of increasing the cutoff radius beyond $7.5\text{\AA}$ (due to the fast increase in the cost of collecting the appropriate training set). For the toluene system, which has a sufficiently complex geometry, a simple increase in the cutoff radius was not enough to achieve the accuracy of the MLIP+D models, especially for the ETN model.
Acknowledgements.
This work was in part supported by the Russian Science Foundation (grant number 23-13-00332, https://rscf.ru/project/23-13-00332/). Max Hodapp acknowledges the financial support under the scope of the COMET program within the K2 Center “Integrated Computational Material, Process and Product Engineering (IC-MPPE)” (Project No 886385); this program is supported by the Austrian Federal Ministries for Climate Action, Environment, Energy, Mobility, Innovation and Technology (BMK) and for Labour and Economy (BMAW), represented by the Austrian Research Promotion Agency (FFG), and the federal states of Styria, Upper Austria and Tyrol.
Author declarations
Conflict of interest
The authors have no conflicts to disclose.
Author Contributions
Olga Chalykh: Conceptualization (equal); Methodology (equal); Software (equal); Formal analysis; Writing - original draft; Writing - review & editing (supporting). Dmitry Korogod: Software (equal); Writing - review & editing (supporting). Ivan S. Novikov: Methodology (equal); Software (supporting); Writing - original draft (supporting); Writing - review & editing (equal). Max Hodapp: Methodology (equal); Software (supporting); Writing - original draft (supporting); Writing - review & editing (supporting). Nikita Rybin: Methodology (supporting); Writing - review & editing (supporting). Alexander V. Shapeev: Conceptualization (equal); Methodology (equal); Writing - review & editing (equal).
Data Availability Statement
The MLIP-2 code used for the fitting and simulations with MTP is publicly available in the GitLab repository at https://gitlab.com/ashapeev/mlip-2.
The MLIP-4 code used for the fitting and simulations with ETN is publicly available in the GitLab repository at https://gitlab.com/ashapeev/mlip-4.