### v2

# Moment Tensor Potential and Equivariant Tensor Network Potential with explicit dispersion interactions

Abstract
In this study, we investigate the effect of incorporating explicit dispersion interactions in the functional form of machine learning interatomic potentials (MLIPs), particularly in the Moment Tensor Potential and Equivariant Tensor Network potential for an accurate modeling of liquid carbon tetrachloride, methane, and toluene. We show that explicit incorporation of dispersion interactions via D2 and D3 corrections significantly improves the accuracy of MLIPs when the cutoff radius is set to the commonly used value of 5–6 Å. We also show that for carbon tetrachloride and methane, a substantial improvement in accuracy can be achieved by extending the cutoff radius to 7.5 Å. However, for accurate modeling of toluene, explicit incorporation of dispersion remains important. Furthermore, we find that MLIPs incorporating dispersion interactions via D2 reach a close level of accuracy to those incorporating D3, implying that D2 is suitable for an accurate modeling of the systems in the study, while being less computationally expensive. We benchmarked the accuracy of the MLIPs on dimer binding curves compared to ab initio data and on predicting density and radial distribution functions compared to experiments.
preprint: AIP/123-QED
I Introduction
Machine learning interatomic potentials (MLIPs) have emerged as widely recognized tools for modeling large systems over extended length- and timescales with the accuracy of ab initio methods Zuo et al. (2020); Kozinsky et al. (2023); Zhang et al. (2025). This makes MLIPs an attractive choice for simulating systems that are, on the one hand, computationally prohibitive for conventional ab initio methods and, on the other hand, beyond the accuracy of (semi-)empirical force fields.
Despite significant advances in the development of MLIPs, several key challenges remain. The primary challenge for MLIPs arises from their locality, which is a consequence of the introduction of a cutoff radius that restricts the number of interacting atoms Zhang et al. (2025). Although this approximation works well for metals and alloys, it requires greater caution for low-dimensional materials, materials with defects, or molecules Babaei and Sadeghi (2024); Behler and Csányi (2021); Anstine and Isayev (2023). Moreover, the locality of MLIPs restricts them from accurately capturing long-range interactions, such as electrostatic, induction, and dispersion Anstine and Isayev (2023), interactions.
MLIPs that explicitly incorporate long-range interactions have already been presented in the literature. Such MLIPs include PhysNet Unke and Meuwly (2019), AIMNet2 Isayev, Anstine, and Zubaiuk (2025), SpookyNet Unke et al. (2021, 2024), TensorMol Yao et al. (2018), and ee4G-HDNNP model Ko et al. (2023), which incorporate both, electrostatic and dispersion interactions. Other MLIPs such as HDNNP potentials Artrith, Morawietz, and Behler (2011); Ko et al. (2021), and the DPLR model Zhang et al. (2022) incorporate electrostatic interactions. At the same time, GAP Bartók et al. (2010) was applied to study systems with strong dispersion interactions Veit et al. (2019); Rowe et al. (2020, 2018); Ibragimova et al. (2025). However, it is also possible to model intramolecular and intermolecular interactions using separate potentials Wengert et al. (2021); Gao and Remsing (2022). Additionally, an alternative approach is to account for long-range interactions via global descriptors Kabylda et al. (2023).
The most common approach to address long-range interactions is to explicitly incorporate physically motivated terms in the total energy expression Anstine and Isayev (2023). Behler proposed a classification of MLIPs into four generations based on how long-range interactions are incorporated Behler (2021); Behler and Csányi (2021). The first and second generations correspond to purely local potentials, while the third generation MLIPs explicitly include long-range interactions, with their strength defined by local features such as atomic charges, dependent on the local environment. The fourth generation accounts for the nonlocality of parameters defining long-range interactions, particularly by making charges dependent on the nonlocal environment. This classification is widely used in the context of MLIPs that incorporate electrostatic interactions, and can also be applied to classify MLIPs that account for dispersion interactions Behler and Csányi (2021).
To explicitly capture dispersion interactions, the most common approach is to utilize one of the dispersion corrections developed for density functional theory (DFT) calculations Grimme et al. (2016); Hermann, DiStasio Jr, and Tkatchenko (2017), such as Tkatchenko-Scheffler (TS) Tkatchenko and Scheffler (2009), exchange-dipole moment (XDM) Becke and Johnson (2005), D3 Grimme et al. (2010), or D2 Grimme (2006) models. For electronic density-dependent dispersion correction schemes, such as TS or XDM, machine learning can be used to predict the most computationally demanding parameters, such as atomic moments and volumes in XDM Tu et al. (2023, 2024). Further optimization of the computational cost can be achieved by using averaged parameters instead of machine learning: atomic volumes for TS Veit et al. (2019); Muhli et al. (2021); Ibragimova et al. (2025) or moments and atomic volumes for XDM Tu et al. (2023). The latter approach has been shown to have an accuracy close to that of the scheme using machine-learned coefficients Tu et al. (2023). The D2 or D3 dispersion corrections, which do not require electronic density for evaluation, can be applied on top of a local potential Wang and Huang (2020); Ying and Fan (2023); Unke and Meuwly (2019); Yao et al. (2018), without machine learning of the dispersion correction parameters. An alternative involves parameterizing the Lennard-Jones potential based on relevant configurations, such as the exfoliation curve for black phosphorus or graphene Rowe et al. (2018); Deringer, Caro, and Csányi (2020).
Another challenge is the efficient generation of high-quality ab initio training sets, which is crucial to ensure accuracy within the training domain, for efficient construction of training set, and to improve the transferability of the resulting potential. To address this challenge, various data-driven approaches have been developed, focusing on optimization of training sets rather than increasing model complexity. These approaches include different variations of active learning (AL), such as methods based on the Query by Committee strategy Smith et al. (2018), Bayesian inference Vandermause et al. (2020), optimization of entropy of atomic descriptors Montes de Oca Zapiain et al. (2022), and the D-optimality criterion employed together with the MaxVol algorithm Podryabinkin and Shapeev (2016).
In the present work, we focus on the development of MLIPs with explicit treatment of the dispersion interaction, particularly Moment Tensor Potentials (MTPs) Shapeev (2016) and Equivariant Tensor Network (ETN) potentials Hodapp and Shapeev (2024). MTP is implemented in the MLIP-2 package, together with the AL approach based on the D-optimality criterion Novikov et al. (2020), which enables efficient training set generation for MTP. ETN is implemented in the MLIP-4 package Shapeev (2024). MTP has been widely applied to the modeling of inorganic materials Korotaev et al. (2019); Podryabinkin et al. (2019); Wang et al. (2020), metals and alloys Marmolejo-Tejada and Mosquera (2022); Tasnadi et al. (2021); Gil and Oda (2022); Zheng et al. (2023), and molten salts Rybin et al. (2024); Sun et al. (2024); Attarian, Morgan, and Szlufarska (2022). However, it was not extensively used to study molecular compounds Rybin, Novikov, and Shapeev (2025); Klimanova, Rybin, and Shapeev (2024); Reicht et al. (2024); Novikov et al. (2024). At the same time, ETN has recently been developed Hodapp and Shapeev (2024) and has been successfully applied to defects in metallic alloys Hodapp (2025), but has not yet been widely benchmarked in various applications.
Here, our objective is to improve the accuracy and applicability of the MTP and ETN models by explicitly incorporating dispersion interactions via D2 and D3 corrections. In our study, we compare the precision of local MLIPS (MLIPS(l)) with MLIPS with explicit dispersion correction (MLIP+D). We study the accuracy differences between MLIP(l)’s and MLIP+D’s when the cutoff radius is set to a value within a commonly used range of 5–6 Å, thereby constraining the MLIP(l)’s ability to capture the full range of dispersion interactions. We additionally test an alternative approach in which we simply extend the cutoff radius for MLIP(l)’s to 7.5 Å, allowing them to describe the major part of dispersion interactions, and compare the accuracy of the resulting MLIP(l)’s with MLIP+D’s.
This comparative analysis was performed on liquid carbon tetrachloride, methane, and toluene simulations, since these systems exhibit predominant dispersion interactions and are widely represented in the literature. Moreover, carbon tetrachloride and toluene are common nonpolar solvents. To examine the predictive power of potentials, we calculate the binding curves and compare them with the ones obtained with ab initio calculations. We also calculate the density and radial distribution function (RDF) with the fitted MLIPs and compare them to the experimental ones.
This study is devoted to benchmarking the above MLIPs fitted to PBE+D2/D3 on the density, binding curves, and RDF, as the aim of this study is to investigate the change in accuracy of MLIPs after incorporating explicit dispersion interactions in their functional form. We do not consider more complex properties, such as diffusion coefficients, whose accuracy may be affected by the level of theory or quantum nuclear effects, as this investigation is beyond the scope of this work.
II Methodology
II.1 Moment Tensor Potential
Moment Tensor Potential (MTP) is a machine learning interatomic potential which was first proposed for single-component materials Shapeev (2016) and generalized to multi-component materials Gubaev, Podryabinkin, and Shapeev (2018). MTP energy $E^{\rm MTP}$ is the sum of contributions $V^{\rm MTP}(\mathfrak{\bm{n}}_{i})$ of atomic neighborhoods ${\bf\mathfrak{n}}_{i}$ for $N$ atoms
$E^{\rm MTP}=\sum\limits_{i=1}^{N}V^{\rm MTP}(\mathfrak{\bm{n}}_{i}).$ (1)
Each neighborhood is a tuple
$\mathfrak{n}_{i}=(\{r_{i1},z_{i},z_{1}\},\ldots,\{r_{ij},z_{i},z_{j}\},\ldots,\{r_{iN_{\rm nbh}},z_{i},z_{N_{\rm nbh}}\}),$
where $r_{ij}$ are relative atomic positions, $z_{i}$, $z_{j}$ are the types of central and neighboring atoms, and $N_{\rm nbh}$ is the number of atoms in neighborhood. Each contribution $V^{\rm MTP}(\mathfrak{\bm{n}}_{i})$ in the potential energy $E^{\rm MTP}$ expands through a set of basis functions
$V^{\rm MTP}({\bf\mathfrak{n}}_{i})=\sum\limits_{\alpha}\xi_{\alpha}B_{\alpha}({\mathfrak{\bm{n}}}_{i}),$ (2)
where $B_{\alpha}$ are the MTP basis functions, $\xi_{\alpha}$ are the linear parameters to be found. To define the functional form of the MTP basis functions we introduce the so-called moment tensor descriptors:
$M_{\mu,\nu}({\mathfrak{\bm{n}}}_{i})=\sum_{j=1}^{N_{\rm nbh}}f_{\mu}(|r_{ij}|,z_{i},z_{j})r_{ij}^{\otimes\nu}.$ (3)
The descriptor consists of the angular part $r_{ij}^{\otimes\nu}$ (the symbol $\otimes$” denotes the outer product of vectors and, thus, the angular part is the tensor of $\nu$-th order) and the radial part $f_{\mu}(|r_{ij}|,z_{i},z_{j})$ of the following form:
$\displaystyle f_{\mu}(|r_{ij}|,z_{i},z_{j})=\sum_{\beta}c^{(\beta)}_{\mu,z_{i},z_{j}}T^{(\beta)}(|r_{ij}|)(R_{\rm cut}-|r_{ij}|)^{2}.$ (4)
Here $\mu$ is the number of the radial function $f_{\mu}$, $c=\{c^{(\beta)}_{\mu,z_{i},z_{j}}\}$ are the radial parameters to be found, $T^{(\beta)}(|r_{ij}|)$ are polynomial functions, and the term $(R_{\rm cut}-|r_{ij}|)^{2}$ is introduced to ensure smoothness with respect to the atoms leaving and entering the sphere with the cutoff radius $R_{\rm cut}$.
By definition, the MTP basis function $B_{\alpha}$ is a contraction of one or more moment tensor descriptors, yielding a scalar. In order to construct the basis functions $B_{\alpha}$, and, thus, determine a particular functional form of MTP, we define the so-called level of moment tensor descriptor:
${\rm lev}M_{\mu,\nu}=2+4\mu+\nu.$ (5)
We also define the level of MTP basis function:
${\rm lev}B_{\alpha}=\rm{lev}\underbrace{\prod_{p=1}^{P}M_{\mu_{p},\nu_{p}}}_{scalar}=\sum\limits_{p=1}^{P}(2+4\mu_{p}+\nu_{p}).$ (6)
A set of MTP basis functions and, thus, a particular functional form of MTP depends on the maximum level, $\rm lev_{\rm max}$}, which we also call the level of MTP. In the set of MTP basis functions, we include only those with ${\rm lev}B_{\alpha}\leq{\rm lev_{\rm max}}$.
II.2 Equivariant Tensor Network Potentials
Polynomial-based potentials, like MTPs, can be generally written as a contraction of a tensor $T$, containing the parameters of the potential, with feature vectors $v$ that contain the same radial and angular information as the moment tensor descriptors (3) (we will make this precise further below)
$V({\bf\mathfrak{n}}_{i})=T_{k_{1}\ldots k_{d}}v^{1}_{k_{1}}\ldots v^{d}_{k_{d}},$ (7)
where the dimension $d$ defines the body-order of the potential. However, Equation (7) is not convenient because the evaluation scales exponentially with $d\cdot n$, where $n$ is the number of features contained in each of the $v$’s. For MTPs, this problem is to some extent solved semi-empirically, e.g., by contracting the radial features before entering the moment tensor descriptors.
A way to explicitly control the growth of the potential’s complexity is to represent $T$ with a tensor network. Tensor networks are a way of representing high-dimensional tensors in a low-rank format by factorizing the full tensor into smaller tensors, up to the order of three. There are different formats, with tensor trains, hierarchical Tucker, or PEPS, arguably being among the most popular ones Cichocki et al. ; Orús , but their contractions with vectors $v^{1}_{k_{1}}$, $v^{2}_{k_{2}}$, etc., can all be realized by sequences of contractions of up-to-order-three tensors
$u^{1}_{k_{1}^{\prime}}$ $=T^{1}_{k_{1}^{\prime}k_{1}}v^{1}_{k_{1}},$
$u^{2}_{k_{2}^{\prime}}$ $=T^{2}_{k_{2}^{\prime}k_{2}k_{1}^{\prime}}v^{2}_{k_{2}}u^{1}_{k_{1}^{\prime}},$
$u^{3}_{k_{3}^{\prime}}$ $=T^{3}_{k_{3}^{\prime}k_{3}k_{2}^{\prime}}v^{3}_{k_{3}}u^{2}_{k_{2}^{\prime}},$
$\ldots.$
The main difference between ETNs and conventional tensor networks is the implementation of symmetry constraints that render the ETN invariant under the actions of the corresponding symmetry group. In our case of interatomic potentials, we require that the energy per atom $V({\bf\mathfrak{n}}_{i})$ remains invariant under the actions of the group of rotations SO(3). [^1] To encode the full O(3) invariance (rotations and reflections) into ETNs, it suffices that $V({\bf\mathfrak{n}}_{i})$ is a real quantity (more details are provided in the original paper Hodapp and Shapeev (2024), section 3.4.2) To that end, we require the feature vectors to be SO(3)-equivariant vectors, that rotate correspondingly with a basis change under SO(3). In the following, we consider a decomposition of each $v$ into an irreducible covariant representation of SO(3) using spherical harmonics. Thus, each index $k$ of a feature vector $v_{k}$ is a multi-index $k=(\ell mn)$, with $\ell=0,\ldots,L$ being the index of the subspace of the irreducible representation, $m\in\{-\ell,-\ell+1,\ldots,\ell\}$ being the dimension of the subspace, and $n=1,\ldots,N(\ell)$ is the number of radial channels corresponding to each $\ell$. We point out that we intentionally deviate from the ordering $n\ell m$, commonly used in quantum physics, that puts the index $n$ in front of $\ell$ and $m$. We chose this notation because each multi-index in our tensor network may depend on a different $\ell$. Therefore, since $m$ and $n$ are always assumed to depend on $\ell$, the ordering $\ell mn$ appears to be more comprehensible in our context.
With this definition of the feature vectors, a sufficient condition for $V({\bf\mathfrak{n}}_{i})$ being invariant under actions of SO(3) is that the tensors $T$ are equivariant maps of these covariant vectors. According to the Wigner-Eckhart Theorem, any $T$ with three multi-indices $\{(\ell_{i},m_{i},n_{i})\}_{i=1,\ldots,3}$, can be factorized to
$T_{(\ell_{1}m_{1}n_{1})(\ell_{2}m_{2}n_{2})(\ell_{3}m_{2}n_{3})}=\theta_{(\ell_{1}n_{1})(\ell_{2}n_{2})(\ell_{3}n_{3})}C_{(\ell_{1}m_{1})(\ell_{2}m_{2})(\ell_{3}m_{3})},$
where $\theta_{(\ell_{1}n_{1})(\ell_{2}n_{2})(\ell_{3}n_{3})}$ is the tensor of model coefficients, and $C_{(\ell_{1}m_{1})(\ell_{2}m_{2})(\ell_{3}m_{3})}$ is the Clebsch-Gordan coefficient that defines the symmetry group. As a tensor network, we use an (equivariant) tensor train representation Oseledets (2011) of atomic energies with equal feature vectors $(v^{1},\ldots,v^{d})=v$. An ETN potential in the tensor train format can then be written as follows
$V^{\rm ETN}({\bf\mathfrak{n}}_{i})=\left(T^{1}_{(\ell_{1}^{\prime}m_{1}^{\prime}n_{1}^{\prime})(\ell_{1}m_{1}n_{1})}v_{(\ell_{1}^{\prime}m_{1}^{\prime}n_{1}^{\prime})}({\bf\mathfrak{n}}_{i})\right)\\ \left(T^{2}_{(\ell_{1}m_{1}n_{1})(\ell_{2}^{\prime}m_{2}^{\prime}n_{2}^{\prime})(\ell_{2}m_{2}n_{2})}v_{(\ell_{2}^{\prime}m_{2}^{\prime}n_{2}^{\prime})}({\bf\mathfrak{n}}_{i})\right)\\ \ldots\left(T^{d}_{(\ell_{d-1}m_{d-1}n_{d-1})(\ell_{d}^{\prime}m_{d}^{\prime}n_{d}^{\prime})}v_{(\ell_{d}^{\prime}m_{d}^{\prime}n_{d}^{\prime})}({\bf\mathfrak{n}}_{i})\right)$ (8)
in which the channels $n_{1},n_{2},\ldots,n_{d}$ naturally emerge as ranks of the tensor network. As feature vectors, we use the ones that contract the radial features before entering the tensor train Hodapp and Shapeev (2024) as follows:
$v_{(\ell mn)}=\sum_{j}\Big{(}B_{\ell n\alpha\lambda}Q_{\alpha}(|r_{ij}|)\big{(}A_{\ell\lambda\beta\gamma}z_{\beta}^{i}z_{\gamma}^{j}\big{)}\Big{)}Y_{\ell m}(r_{ij}/|r_{ij}|).$
When comparing the moment tensor descriptors $M_{\mu,\nu}({\mathfrak{\bm{n}}}_{i})$ and the feature vectors of ETNs, the radial parameters $c$ are decomposed into $A$ and $B$, and the angular part $r_{ij}^{\otimes\nu}$ is reshaped to $Y_{\ell m}(r_{ij}/|r_{ij}|)$. More details on the relation between MTP and ETN are provided in the original paper by Hodapp and Shapeev Hodapp and Shapeev (2024).
This enables learning similarities between the all radial features from the data through the parameter tensors $A$ and $B$, avoiding the problem of an exponentially growing size of the feature vectors Hodapp and Shapeev (2024). As for MTP, the total energy of ETN is the sum of contributions $V^{\rm ETN}({\bf\mathfrak{n}}_{i})$ for $N$ atoms.
The number of coefficients of the ETN potentials is proportional to $d\bar{r}^{2}\bar{n}$, where $\bar{r}$ is the average rank of the tensor network, and $\bar{n}$ is the average size across all dimensions of the feature vector. This makes the growth of their parameter space more controllable than, e.g., for MTPs.
II.3 Fitting
For finding parameters of MLIPs, e.g. MTP or ETN parameters, we fit it on a (quantum-mechanical) training set. Let $K$ be a number of configurations in the training set and $N$ be a number of atoms. Denote a vector of MTP/ETN parameters to be found by $\bm{\theta}$}. We find the optimal parameters $\bm{\bar{}}{\theta}$} by solving the following optimization problem (minimization of the objective function)
$\begin{array}[]{c}\displaystyle\sum\limits_{k=1}^{K}\Bigl{[}w_{\rm e}\left(E^{\rm MLIP}_{k}({\bm{\theta}})-E^{\rm QM}_{k}\right)^{2}+\\ \displaystyle w_{\rm f}\sum_{i=1}^{N}\left|{\bf f}^{\rm MLIP}_{i,k}({\bm{\theta}})-{\bf f}^{\rm QM}_{i,k}\right|{{}^{2}}\Bigr{]}\to\operatorname{min}.\end{array}$ (9)
We start from randomly initialized MLIP parameters. The optimal parameters $\bm{\bar{}}{\theta}$} are found numerically, using the iterative method to minimize the non-linear objective function, namely the Broyden-Fletcher-Goldfarb-Shanno algorithm. Thus, after optimization, the parameters $\bm{\bar{}}{\theta}$} are near the local minimum of the objective function. In this objective function, $E^{\rm QM}_{k}$ and ${\bf f}^{\rm QM}_{i,k}$ are the reference energies and forces, i.e., the ones to which we fit the MLIP energies $E^{\rm MLIP}_{k}$ and forces ${\bf f}^{\rm MLIP}_{i,k}$ and thus optimize the MLIP parameters $\bm{\theta}$}. The factors $w_{\rm e}$ and $w_{\rm f}$ are non-negative weights which express the importance of energies and forces with respect to each other. We took $w_{\rm e}=1$ and $w_{\rm f}=0.01$ to fit the ETN-based models and we took $w_{\rm e}=1/N$ and $w_{\rm f}=0.01$ to train the MTP-based models. These weights are default in the MLIP-2 and MLIP-4 codes, respectively. We refer to the minimization problem (9) as MLIP fitting.
II.4 Active Learning
Assume that we found the vector of the optimal MTP parameters $\bm{\bar{}}{\theta}$} after solving (9) and let the length of the vector be $m$. We then construct a matrix
$\mathsf{B}=\left(\begin{matrix}\frac{\partial E^{\rm MTP}_{1}}{\partial\theta_{1}}({\bm{\bar{\theta}}})&\ldots&\frac{\partial E^{\rm MTP}_{1}}{\partial\theta_{m}}({\bm{\bar{\theta}}})\\ \vdots&&\vdots\\ \frac{\partial E^{\rm MTP}_{K}}{\partial\theta_{1}}({\bm{\bar{\theta}}})&\ldots&\frac{\partial E^{\rm MTP}_{K}}{\partial\theta_{m}}({\bm{\bar{\theta}}})\end{matrix}\right).$
From this matrix, we select a set of the $m$ most linearly independent rows, i.e., we find a submatrix $A$ with the maximum absolute value of the determinant (maximum volume). To find a submatrix $A$ of maximum volume from the matrix $B$, we use the MaxVol algorithm Goreinov et al. (2010) which ensures the diversity of the configurations corresponding to the rows of $A$.
After training the initial MTP and constructing the matrix $A$, we start a molecular dynamics (MD) simulation and compute the extrapolation grade for each configuration $\bm{x}^{*}$} created during this simulation
$\gamma({\bm{x}^{*}})=\max\limits_{1\leq j\leq m}|c_{j}|,$ (10)
where the vector $\textbf{c}=(c_{1},\ldots,c_{m})$ for the configuration $\bm{x}^{*}$} is computed by
$\textbf{c}=\Bigl{(}\frac{\partial E^{\rm MTP}}{\partial\theta_{1}}(\bm{\bar{\theta}},{\bm{x}}^{*}),\dots,\frac{\partial E^{\rm MTP}}{\partial\theta_{m}}(\bm{\bar{\theta}},{\bm{x}}^{*})\Bigr{)}A^{-1}.$ (11)
We also introduce two thresholds: $\gamma_{\rm save}$ and $\gamma_{\rm break}$. Now we have all the prerequisites to formulate our AL algorithm.
  1. 1.
    Run the MD simulation with extrapolation control. The MD simulation algorithm generates configurations for which energy and forces are computed. An extrapolation grade is also calculated for each configuration. Depending on the degree of extrapolation $\gamma({\bm{x}^{*}})$ there are three possibilities: (i) if $\gamma({\bm{x}^{*}})<\gamma_{\rm save}\approx 2$ then we proceed with the calculation of the energy and forces as we deal with interpolation or insignificant extrapolation; (ii) if $\gamma_{\rm save}<\gamma({\bm{x}^{*}})<\gamma_{\rm break}\approx 10$ then we save the configuration to a file with the extrapolative configurations and proceed with the calculation of the energy and forces; (iii) if $\gamma({\bm{x}^{*}})>\gamma_{\rm break}$ then we break the calculation as the extrapolation degree exceeds the critical threshold $\gamma_{\rm break}$.
  2. 2.
    Selection of configurations. In the previous step, the extrapolative configurations were saved to the file. However, the saved file may contain a large number of similar configurations. According to our active learning strategy, we select among the extrapolative configurations those that maximize the volume of the matrix $A$. Thus, in this step, we select the configurations to be added to the training set.
  3. 3.
    Quantum-mechanical calculations. After selecting the appropriate configurations, we calculate their quantum-mechanical energy and forces.
  4. 4.
    MTP retraining. After ab intio calculations of the configurations, we add them to the training set, retrain MTP, and update the matrix $A$.
We repeat steps 1-4 until no configuration is preselected during MD simulations. We note that the active learning algorithm can only be used with MTP and was implemented only in the MLIP-2 code Novikov et al. (2020). For the moment, we have not implemented a similar active learning algorithm for ETN.
II.5 Range separation
In the present study, we added the D2 Grimme (2006) or D3 Grimme et al. (2010) dispersion corrections to the local MTP and ETN potentials. We implemented the D2 correction in both the MLIP-2 and MLIP-4 codes, adding it to both MTP and ETN. The D3 correction was implemented only in the MLIP-4 code and combined only with ETN.
The local potentials, namely MTP and ETN, were trained on configurations with total energies and forces. In contrast, potentials that explicitly incorporate dispersion interactions via D2 or D3, namely, MTP+D2, ETN+D2, ETN+D3, were obtained in the following way. We first trained pure MTP and ETN models on configurations with D2 or D3 contributions subtracted from total energies and forces, and after training, dispersion correction was explicitly added to the top of a trained MLIP.
The explicit incorporation of dispersion via D2 or D3 allows to maintain a cutoff radius of 5–6 Å for the machine learning part of these potentials, while the cutoff radius for D2 or D3 is set to 15 Å for all simulations presented in this work.
II.6 D2 correction
The D2 dispersion correction was introduced by Grimme et al. Grimme (2006) to account for dispersion interactions that are not inherently included in the DFT framework.
D2 considers only the two-body dipole-dipole contribution to the dispersion energy, leading to the following form of energy expression:
$E^{D2}=-s_{6}\sum_{ij}\frac{C_{6}^{ij}}{r_{ij}^{6}}f_{\rm damp}(r_{ij}),$ (12)
where $i,j$ are indices of atoms, $s_{6}$ is a scaling factor dependent on a particular density functional, $C_{6}^{ij}$ is a coefficient describing interactions between atoms $i$ and $j$, $r_{ij}$ is an interatomic distance between atoms $i$ and $j$, and $f_{\rm damp}(r_{ij})$ is a damping function.
In this scheme, the coefficients $C_{6}^{ii}$ are precomputed parameters and are calculated from the empirical formula:
$C_{6}^{ii}=0.05NI_{A}\alpha_{A}^{0},$ (13)
where $N$ is a scaling factor, $I_{A}$ is the first ionization potential calculated with PBE0, and $\alpha_{A}^{0}$ is static dipole polarizability. The $C_{6}^{ij}$ are defined as the geometric mean of $C_{6}^{ii}$ and $C_{6}^{jj}$:
$C_{6}^{ij}=\sqrt{C_{6}^{ii}C_{6}^{jj}}.$ (14)
The Fermi damping function $f^{\rm Fermi}_{\rm damp}$ is utilized in D2:
$f^{\rm Fermi}_{\rm damp}=\frac{1}{1+e^{{20(r_{ij}/(s_{R}R_{\rm vdW})-1)}}},$ (15)
where $s_{R}$ is a scaling parameter, $R_{\rm vdW}$ is a sum of van der Waals radii of interacting atoms.
However, in our implementation of the D2 correction (as well as the D3 correction), we chose Becke-Johnson damping function $f^{(n)}_{\rm damp,BJ}$:
$f^{(n)}_{\rm damp,BJ}=\frac{r^{n}}{r^{n}+(a_{1}R_{0}+a_{2})^{n}}$ (16)
with $n=6$. We adjusted the coefficients $a_{1}$, $a_{2}$, and $R_{0}$ on CCl4, CH4, and toluene dimer binding curves.
II.7 D3 correction
In contrast to the D2 scheme, the D3 dispersion correction Grimme et al. (2010) explicitly takes into account information about the atomic environment via fractional coordination number. Moreover, in the D3 approach, the dispersion energy includes not only the dipole-dipole interaction term, proportional to $1/r^{6}$, but also the higher order dipole-quadrupole contribution proportional to $1/r^{8}$:
$E^{D3}=-\sum_{ij}\sum_{n=6,8}s_{n}\frac{C_{n}^{ij}}{r_{ij}^{n}}f_{\rm damp}^{(n)}(r_{ij}),$ (17)
where $i,j$ are the indices of atoms, $s_{n}$ is a scaling factor dependent on the density functional, $C_{n}^{ij}$ is a coefficient describing interactions between atoms $i$ and $j$, $r_{ij}$ is the interatomic distance between atoms $i$ and $j$, and $f_{\rm damp}^{(n)}(r_{ij})$ is damping function, dependent on the power $n$.
It is possible to include the three-body interaction term in the expression of the D3 energy. However, in our implementation, we limited D3 to the two-body form, as this implementation is widely used and the contribution of the three-body interaction term is typically much lower than the two-body term Grimme et al. (2016).
In the D3 correction, $C_{n}^{ij}$ explicitly includes information on the environments of both $i$ and $j$ atoms via coordination numbers:
$CN^{i}=\sum_{i\neq j}^{N_{\rm atoms}}\frac{1}{1+e^{{-16(4(R_{i,{\rm cov}}+R_{j,{\rm cov}})/(3r_{ij})-1)}}},$ (18)
where $R_{i,{\rm cov}}$, $R_{j,{\rm cov}}$ are covalent radii of elements Pyykkö and Atsumi (2009).
The computed coordination numbers are further used for interpolating $C_{6}^{ij}$ based on a set of precomputed coefficients $C_{6,\rm ref}^{ij}$, obtained for reference compounds Grimme et al. (2010):
$C_{6}^{ij}(CN^{i},CN^{j})=\frac{\sum_{p}^{N_{i}}\sum_{q}^{N_{j}}C_{6,\rm ref}^{ij}(CN_{p}^{i},CN_{q}^{j})L_{pq}}{\sum_{p}^{N_{i}}\sum_{q}^{N_{j}}L_{pq}},$ (19)
where $N_{i}$, $N_{j}$ are numbers of reference compounds for elements $i$ and $j$, and $L_{pq}$ is defined as:
$L_{pq}=e^{{-4\big{(}(CN^{i}-CN^{i}_{p})^{2}+(CN^{j}-CN^{j}_{q})^{2}\big{)}}}.$ (20)
Based on the obtained $C_{6}^{ij}$, the coefficients $C_{8}^{ij}$ for the higher-order term in dispersion energy (17) is computed:
$C_{8}^{ij}=3C_{6}^{ij}\sqrt{Q_{i}Q_{j}},$ (21)
where $Q_{i}$, $Q_{j}$ are precomputed parameters derived from atomic densities and atomic numbers of elements $i$ and $j$ Grimme et al. (2010).
For the D3 correction, two types of damping function can be utilized: zero-damping and Becke-Johnson damping (BJ) (16). For our implementation of D3 we chose the BJ damping with $n=6$ for the $1/r^{6}$ damping term, and $n=8$ for the higher order term.
II.8 Protocol of datasets generation and training of MLIPs
We collected training sets with total energies and forces via the AL approach described above and implemented in the MLIP-2 code for MTP. For fitting of models that explicitly incorporate dispersion via D2 or D3 (MTP+D2, ETN+D2, ETN+D3), the training set with total energies and forces was post-processed by subtracting D2 or D3 contributions from the total energies and forces. The general workflow for obtaining training sets for each of the considered models is shown in Figure 1.
Figure 1: The general scheme for generating training sets for all the MLIPs considered. At the first step, the AIMD simulation is conducted and initial MTP is trained on the configurations uniformly chosen from the AIMD trajectory. Next, a pretraining set formed from uniformly selected configurations and configurations, additionally selected from the AIMD trajectories with the MaxVol algorithm Podryabinkin and Shapeev (2016). The initial MTP retrained on the resulting pretraining set. After that, the AL of MTP during MD simulations is carried out, and results in the actively selected training set with total energies and forces. Further steps correspond to training of MTP and ETN on the dataset including total energies and forces, and subtracting D2 or D3 contributions and utilizing the resulting training sets for fitting of MTP or ETN part of MTP+D2, ETN+D2, ETN+D3 potentials.
We started creating a training set from ab initio molecular dynamics (AIMD) simulations. An initial configuration was generated using Packmol Mart´ınez et al. (2009). Next, we uniformly chose about 20 configurations from this AIMD trajectory to avoid correlations between configurations and fit an initial MTP. After that, we selected configurations using the MaxVol algorithm from the pool of configurations generated during AIMD, updated the initial training set, and refitted the MTP, obtaining pretrained MTP. Subsequently, we actively trained the pretrained MTP during 10 parallel MD simulations (AL-MD) of 100 ps each using the AL procedure described above, utilizing LAMMPS package Thompson et al. (2022) for MD. As a result, we obtained the actively selected training set with total energies and forces. During both AIMD and MD in the AL procedure, a timestep was set to 1 fs.
The collected training set was then used to train the MTP and ETN models. After subtracting the contributions of D2 or D3 from the total energies and forces in the collected training set, we obtained training sets for the MTP+D2, ETN+D2, and ETN+D3 models.
For all the investigated systems, the level of the MTP-based potentials (MTP and MTP+D2) was set to 16, corresponding to 222 parameters to be optimized during the fitting, and for the ETN-based models (ETN, ETN+D2, and ETN+D3) we took 124 coefficients to be fitted. We note that the AL procedure was performed only for MTP and the resulting training set may not be optimal for other tested MLIPs, especially for the ETN-based models. We emphasize that in this procedure we implicitly assume, since ETN has fewer parameters, that the training set actively selected by MTP would also be suitable for ETN.
For creating validation sets, the configurations were uniformly sampled from independent 100 ps MD simulation conducted with actively trained MTP. The energies and forces for the sampled configurations were then computed using ab initio calculations. The ratio between the size of the training and validation sets was set to 8:2.
II.9 Assessing the quality of MLIPs
To compare the quality of the fitted MLIPs we evaluated training and validation errors of predicting energies and forces. In addition, we designed two series of tests. In the first series, we evaluated the accuracy of potentials in predicting intermolecular interactions. To that end, we predicted binding curves for a diverse set of dimers and compared them with the ones obtained with ab initio calculations. We calculated the deviation between the position and depth of the minima, as well as the full width of the well at half its depth (referred to as full-width at half-minimum or FWHM) obtained with potentials and ab initio calculations (see supplementary material for details). In this test, the binding curves were represented as dependencies of the interaction energy on the distance between the centers of mass of individual molecules. The second series of tests aimed to assess the accuracy of density and RDF calculated during MD, performed with LAMMPS package Thompson et al. (2022). We used experimental data as reference.
For evaluating the errors, we used the root mean square error (RMSE) and mean absolute percentage error (MAPE):
${\rm RMSE}=\sqrt{\frac{\sum_{i=1}^{N}(x_{i}-\hat{x}_{i})^{2}}{N}},$ (22)
${\rm MAPE}=\frac{\sum_{i=1}^{N}|x_{i}-\hat{x}_{i}|}{N},$ (23)
where $N$ is the number of samples, and $x_{i}$ and $\hat{x}_{i}$ are the predicted and true values, respectively.
To ensure that error estimations are statistically meaningful, we conducted each test for an ensemble of five potentials for each type of MLIP model considered, except for calculating density and RDF with ETN+D3, as this type of MLIP is more computationally demanding than others.
III Results and discussion
In this section, we present results of application of the described MLIPs to the modeling of liquid carbon tetrachloride, liquid methane, and liquid toluene. For all systems, we focus on evaluating the accuracy differences between MLIP(l)’s and MLIP+D’s.
We note that for carbon tetrachloride, we performed two test cases. In the first case, we compared the accuracy of MLIP(l)’s and MLIP+D’s when the $R_{\rm cut}$ for all potentials is set to 5.5 Å, which limits the ability of MLIP(l)’s to fully capture the range of dispersion interactions. In the second case, we explored to what extent the accuracy of MLIP(l)’s can be improved by increasing the $R_{\rm cut}$ to 7.5 Å, allowing them to capture the major part of long-range interactions (while maintaining a cutoff radius of 5–6 Å for MLIP+D’s). For methane and toluene, we took only the $R_{\rm cut}$ of 7.5 Å for the MLIP(l)’s and of 6 Å for MLIP+D’s.
III.1 Computational details
For the generation of the training set, we employed the AL scheme described above (Figure 1). The simulation setup for AIMD and AL-MD, details of the ab initio calculations, and sizes of the obtained training sets are presented in Table 1.
ab initio package level of theory basis set pseudo-potential cutoff,
k-points
number of
molecules
in cell
cell size, Å MD training
set size
[…]
QZV2P-GTH-q4,
QZV2P-GTH-q7
GTH-PBE-q4,
GTH-PBE-q7
energy cutoff VandeVondele et al. (2005): 400 Ry; $\Gamma$-point 6 10-12 AIMD: at 298 K for 1 ps in the NVT ensemble;
AL-MD: NpT ensemble at 293 K and 1 bar
478
[…]
$\Gamma$-point
11 15-17 AIMD: 500 ps was carried out in the NVE, velocities initialized at 298 K;
AL-MD: NpT ensemble at 293 K and 1 bar
850
CH4 VASP PBE-D3 PW PAW_PBE 900 eV
$\Gamma$-point
57 15-18 AIMD: NVE ensemble for 500 fs, velocities initialized at 298 K;
AL-MD: NpT ensemble at 150 K and 10 bar.
306
Toluene FHI-aims Blum et al. (2009) PBE-TS Numerically tabulated Atom-centered Orbitals Blum et al. (2009), intermediate basis set all electrons 2x2x2 k-point grid 13 15-18 AIMD: NVT ensemble for 400 fs, at 298 K;
AL-MD: NpT ensemble at 298 K and 1 atm
520
Table 1: Computational details on the generation of the training sets.
In all ab initio calculations, we employed dispersion-corrected DFT at the PBE-D3 or PBE-TS level of theory. The employed D3 version includes the Becke-Johnson damping function and omits the three-body (ATM) term. This means that we do not account for many-body effects, which affect the accuracy of the ab initio method and, consequently, the trained potentials. However, since dispersion in ab initio methods is described only by two-body potentials, this approach is consistent with our use of two-body D2 and D3 terms in the MLIP+D framework. However, if MLIP+Dn models were fitted to reference data that included many-body effects, incorporating the ATM term in the D3 term should be expected to enhance the accuracy.
The chosen ab initio models described above were also employed for the calculation of the dimer binding curves, used to test the resulting MLIPs. In addition to testing MLIPs on the binding curves, we also used them during MD simulations to predict densities and RDFs. To determine the densities, we equilibrated the simulation cells in the NpT ensemble until convergence was achieved, and then averaged the density over 100 ps of the simulation. The simulation details of the density calculations are given in Table 2. For calculating RDFs, the simulations started from the configuration with converged density. Then, the system was equilibrated in the NVT ensemble at the target temperature during 10 ps. Subsequently, the ensemble was changed to NVE, and the RDF was averaged over 100 ps. A timestep of 1 fs was used in all simulations. For CCl4 and toluene, the RDF was obtained under the same conditions as the density, while for CH4 the temperature was set to 92 K and the pressure was set to 0.13 atm to match the conditions of the available experimental data.
number of molecules Density simulation
trajectory length, ps T, K p, atm
[…]
CH4 456 300 110 20.48
Toluene 104 200 298 1
Table 2: Computational details on density calculations.
III.2 Training and validation errors
For all the test cases, ensembles of five potentials were trained for each MLIP type. The training and validation RMSEs, averaged over the ensembles, are presented in Figure 2. For MLIPs with a cutoff of 5.5 Å, we observe a reduction in the energy RMSEs moving from MTP to MTP+D2, while the effect of dispersion correction for ETN-based potentials is less significant and observed only for validation energy RMSEs. When extending the cutoff radius of MLIP(l)’s from 5.5 Å to 7.5 Å, the errors decreased by approximately a factor of three for energy, matching the accuracy of MLIP+D’s. This trend also holds for CH4, where errors are approximately four times smaller for energies and two times lower for forces than those obtained for CCl4 with MLIPs with extended $R_{\rm cut}$. However, the MLIPs fitted on toluene, even when the cutoff radius of MLIP(l)’s is extended, were not able to reach the accuracy level of MLIP+D’s.
[…]
III.3 Binding curves
[…]
Figure 3: a) Dimers of CCl4 and CH4 used for generation of the binding curves. The labels correspond to description of mutual orientation, based on treatment of CCl4 molecule as tetrahedron. F corresponds to “face”, V – “vertex”, E – “edge”, and, finally, s and e correspond to staggered and eclipsed conformations; b) dimers of toluene used for constructing binding curves and minima depths and positions. The labels denote mutual orientations of molecules: "P" – parallel, "T" – T-shaped, "IP" – in plane.
The minima depth and position on the binding curves for CCl4, CH4 and toluene dimers obtained with ab initio calculations are presented in Table 3. The binding curves were calculated with the same ab initio package that was used to generate the corresponding training set. Compared to carbon tetrachloride, CH4 dimers exhibit closer and shallower minima. In contrast, toluene dimers possess the widest range of minima depths and positions among the studied systems, suggesting that toluene will be the most challenging molecule to model for MLIPs.
$F$s FFe FVs FVe $E$s EEe EF EV $V$s VVe
[…]
r 4.9 5.2 5.9 5.9 6.0 5.5 5.4 6.2 7.3 7.3
[…]
r 4.9 5.2 5.9 5.9 6.0 5.5 5.4 6.2 7.2 7.2
CH4 E -35 -33 -29 -29 -25 -29 -31 -25 -17 -18
r 3.7 3.8 4.1 4.1 4.0 4.0 3.9 4.2 4.6 4.6
Toluene P-1 P-2 P-3 P-4 T-1 T-2 T-3 IP-1 IP-2 IP-3
E -173 -172 -134 -115 -123 -139 -104 -26 -46 -57
r 3.8 4.2 5.1 5.8 6.0 5.6 5.9 8.2 8.3 8.4
[…]
The calculated RMSEs for the energy minimum ($E_{\rm min}$), the position of its minimum ($r_{\rm min}$), and the FWHM are shown in Figure 4 (a). We observed that explicit incorporation of the dispersion correction significantly affects the FWHM RMSEs when $R_{\rm cut}$ is set to 5.5 Å for MLIP(l)’s, and it also significantly improves the accuracy of $E_{\rm min}$ for the ETN model (Figure 4 (a)). The effect of incorporating the dispersion correction corresponds to approximately 50% decrease in the FWHM RMSE for MTPs-based potentials and to 50% decrease in the energy RMSE for the ETN-based potentials. However, RMSE changes of the same order of magnitude are observed when extending the MLIP(l)’s cutoff radius to 7.5 Å (Figure 4 (a)). In particular, the ETN energy RMSE and MTP FWHM RMSE decrease by about 50%, reaching the accuracy of MLIP+D’s. Accordingly, for the other systems, since the MLIP(l)’s cutoff radii were extended, the difference between MLIP(l)’s and MLIP+D’s is not significant in any property excluding FWHM in CH4 where MLIP+D’s still show higher accuracy than MLIP(l)’s. At the same time, the RMSEs for CH4 are much smaller than for CCl4, while for toluene they are significantly higher than in other test cases, including CCl4 modeling with MLIPs when $R_{\rm{cut}}$ is set to 5.5 Å. Additionally, we note that predicting the depth and FWHM is more challenging than predicting the position of the minimum, as it is less sensitive to the incorporation of dispersion correction or increasing $R_{\rm cut}$.
[…]
Since we observed the significant impact of dispersion correction in modeling of CCl4 with MLIPs when $R_{\rm cut}$ is set to 5.5 Å, we quantified this impact in absolute values of the energy minima and FWHMs (see supplementary material for absolute values of the minima energy and FWHM for other test cases). As shown in Figure 4 (b), the high energy RMSE of ETN corresponds to a significant overbinding of the majority of dimers, which was mitigated by the explicit incorporation of the dispersion correction in ETN+D2 and ETN+D3. At the same time, the difference between the energies predicted by MTP and MTP+D2 is not significant. We also note that all the potentials, excluding ETN+D3, tend to underestimate FWHM, especially the MTP and ETN potentials.
Figure 5: Toluene dimers binding curves predicted by MLIPs compared to ab initio calculated.
In addition to the findings related to the energy minima and the potential well width, we made a further observation concerning the energy dependence on the intermolecular separation. In Figure 5 we demonstrate the unphysical oscillations in the toluene dimer binding curves predicted by MLIP(l)’s with $R_{\rm cut}=7.5$ Å and MLIP+D’s with $R_{\rm cut}=6$ Å. This behavior is observed for all the MLIPs, but is more severe in the case of the ETN-based ones. These oscillations are weaker for MLIP+D’s compared to MLIP(l)’s. This observation explains the difference in MLIP(l)’s and MLIP+D’s training and validation energy errors, which could not be explained from the energy minima, position, and FHWM RMSEs, as they did not change substantially with the incorporation of dispersion correction. Similarly, such oscillations are observed for CCl4 VVe dimer binding curve predicted by MTP with the extended $R_{\rm cut}$ (see supplementary material for details), while MTP+D2 preserved more correct asymptotics. We note that all binding curves for CCl4 and CH4 are also given in supplementary material.
Finally, the overall improvement in binding curve predictions achieved by incorporation of long-range interactions is both intuitive and well-documented in literature for MLIPs that account for electrostatic interactions Grisafi, Nigam, and Ceriotti (2021); Yue et al. (2021).
III.4 Densities
Next, we calculated the liquid densities for all considered cases. Similar to the binding curve tests, we focus on the impact of adding the dispersion correction and, for CCl4, on extending the cutoff radius for MLIP(l)’s.
All density values were obtained by averaging over 100 ps of the NpT MD trajectories after the density converged. The details of the simulation are given in Table 2. Figure 6 shows the predicted and experimental densities, and Table 4 contains the corresponding MAPEs for all investigated systems. Similarly to the binding curve tests, only CCl4 was modeled with MLIP(l)’s and MLIP+D’s using two sets of cutoff radii.
[…]
MTP MTP+D2 ETN ETN+D2 ETN+D3
CCl4, (a) 0.991 0.26 0.26 0.216 0.22
$\pm$ 0.001 $\pm$ 0.04 $\pm$ 0.04 $\pm$ 0.009 $\pm$ 0.02
CCl4, (b) 0.156 0.20 0.170 0.18 -
$\pm$ 0.004 $\pm$ 0.02 $\pm$0.008 $\pm$ 0.01
CH4 0.033 0.045 0.035 0.043 -
$\pm$ 0.004 $\pm$ 0.005 $\pm$ 0.003 $\pm$ 0.002
Toluene 0.12 0.100 0.21 0.14 -
$\pm$ 0.02 $\pm$ 0.006 $\pm$ 0.06 $\pm$ 0.01
[…]
The two key findings during the modeling of the CCl4 system using all MLIPs with cutoff 5.5 Åare given below. First, the MTP potential does not maintain stable molecular dynamics in the liquid phase, as we observed the density collapsing to near zero values. However, this issue is resolved by explicit incorporation of dispersion interactions in MTP+D2, resulting in a 25.6% deviation from the experimentally measured density Hien, Ponter, and Peier (1978) of 1.594 g/cm3 (Figure 6 (a), Table 4). Unlike MTP, the ETN potential is capable of preserving the liquid phase, likely due to the overbinding of molecules, which was demonstrated in the dimer binding curves test (Figure 4 (b)). Additionally, ETN+D2 and ETN+D3 have a higher accuracy compared to ETN, achieving a deviation of approximately 22% from the experimental density. Secondly, unlike the results of the binding curves tests, ETN+D3 does not outperform ETN+D2 in predicting density.
When the cutoff radii were extended, we obtained a substantial improvement over the MLIPs with smaller cutoff radii, as can be seen from the density values and MAPEs (Figure 6 (b), Table 4). Furthermore, similarly to the binding curves, the errors of all MLIPs are close to each other. Thus, we demonstrated that it is sufficient to introduce an $R_{\rm cut}$ capturing the major part of long-range interactions for MLIP(l)’s for accurate modeling of liquid carbon tetrachloride, since the difference in accuracy between the MLIP(l) and MLIP+D potentials is minor for all properties calculated.
In view of all the results for CCl4, for the other two systems, we decided to limit the tests for MLIP+D’s with $R_{\rm cut}=6$ Å and MLIP(l)’s with $R_{\rm cut}=7.5$ Å.
The density of the CH4 system was compared to the experimental value Harris and Trappeniers (1980) of 0.4256 g/cm3 (Figure 6 (c), Table 4). The errors obtained for all MLIPs are small and close in magnitude. Furthermore, no significant differences in the accuracy of MLIP(l)’s and MLIP+D’s are observed similarly to the modeling of CCl4. In contrast, the density of toluene predicted by ETN exhibits a large uncertainty, indicating substantial deviations in the MLIP training results (Figure 6 (d), Table 4). Furthermore, the mean density evaluated with ETN significantly overestimates the experimental value Harris (2000) of 0.862 g/cm3. For the ETN + D2 potential, these issues are resolved, and a density value closer to the experiment is achieved together with a lower spread. MTP also shows higher error and uncertainty compared to MTP+D2, although the difference is not as dramatic as for ETN and ETN+D2.
To sum up, we observed that the MLIP+D’s were more accurate for density calculations than MLIP(l)’s for CCl4 when the cutoff radii were set to 5.5 Å. However, when the cutoff radii of MLIP(l)’s were extended, their accuracy in modeling CCl4 and CH4 improved to the level of MLIP+D’s and even surpassed MTP+D2 for MTP potentials. In contrast, for toluene, ETN+D2 significantly outperformed ETN even when its cutoff radius was extended. This is likely attributed to the unphysical oscillations observed in toluene binding curves (Figure 5), which were significantly reduced after incorporating the dispersion correction.
III.5 Radial distribution functions
Finally, we calculated RDFs, which were obtained by averaging over 100 ps of the NVE MD trajectories after the equilibration of the density in the NpT ensemble and the temperature in the NVT ensemble. The details of the simulations can be found in Section III.1. The RDFs calculated with MLIPs are shown in Figure 7 and Table 5 contains the MAPEs for extrema positions and intensities, compared to those obtained experimentally for CCl4 Lowden and Chandler (1974), CH4 Oobatake, Hayashi, and Machida (1990), and toluene Falkowska et al. (2016).
Figure 7: Radial distribution functions obtained from experiment Lowden and Chandler (1974); Oobatake, Hayashi, and Machida (1990); Falkowska et al. (2016) and with MLIPs for: a) CCl4 using MLIPs with cutoff 5.5 Å; b) CCl4 using MLIP(l)’s with cutoff 7.5 Å and MLIP+D’s with cutoff 6 Å; c) CH4 using MLIP(l)’s with cutoff 7.5 Å and MLIP+D’s with cutoff 6 Å; d) toluene using MLIP(l)’s with cutoff 7.5 Å and MLIP+D’s with cutoff 6 Å.
Error MTP MTP+D2 ETN ETN+D2 ETN+D3
[…]
$\pm$ 0.03 $\pm$ 0.04 $\pm$ 0.05 $\pm$ 0.05
MAPE(r) - 0.04 0.03 0.03 0.03
$\pm$ 0.02 $\pm$ 0.02 $\pm$ 0.02 $\pm$ 0.02
[…]
$\pm$ 0.08 $\pm$ 0.05 $\pm$ 0.09 $\pm$ 0.09
MAPE(r) 0.03 0.03 0.03 0.03 -
$\pm$ 0.01 $\pm$ 0.02 $\pm$ 0.01 $\pm$ 0.01
CH4 MAPE(g) 0.04 0.03 0.04 0.03 -
$\pm$ 0.02 $\pm$ 0.01 $\pm$ 0.01 $\pm$ 0.01
MAPE(r) 0.02 0.02 0.02 0.03 -
$\pm$ 0.02 $\pm$ 0.01 $\pm$ 0.02 $\pm$ 0.02
[…]
All MLIPs with $R_{\rm cut}=5.5$ Å give nearly equivalent RDFs for CCl4 (Figure 7 (a), Table 5), at the same time exhibiting a shift compared to the experimental data due to an underestimation of the density. In contrast to the results obtained for the density, we do not observe a substantial improvement in accuracy after incorporating dispersion corrections in the present case. Note that the RDF was not calculated with MTP as it failed to maintain the liquid phase in the MD simulations. When CCl4 is modeled with MLIP(l)’s with a cutoff of 7.5 Å, and MLIP+D’s with cutoff 6 Å, the predictions of all MLIPs are close to each other (Figure 7 (b), Table 5), similarly to the calculations using MLIPs with smaller cutoffs.
For the CH4 system, similarly to the accurate density predictions, the resulting RDFs are close to the experimental ones and are nearly the same for all (see Figure 7 (c), Table 5). Additionally, we note that the pressure conditions of the simulations for density and RDF calculations differ significantly from those during the MD simulations in the AL process, yet the results are highly accurate, indicating excellent transferability of the trained MLIPs.
For modeling toluene, the configurations were collected every 50 fs and post-processed using MDAnalysis Michaud-Agrawal et al. (2011). The RDFs were calculated based on the C4 carbon atoms (opposite to the methyl group) and the C7 carbon atoms (methyl carbon). Three types of RDFs were computed: C7–C4, C4–C4, and C7–C7 (Figure 7 (d)). Note that in the C7–C4 RDF the intramolecular contribution was subtracted.
RDFs obtained with all the MLIPs, except for ETN, show qualitative agreement with the experiment. In contrast, the average C4–C4 RDF predicted with ETN has a false maximum. In addition, individual RDFs predicted by individual ETN potentials in the ensemble demonstrate significant spread from each other, leading to broad uncertainty regions. For the ETN+D2 potential, the width of the uncertainty region is substantially reduced and a better agreement with the experiment is achieved. The difference between MTP and MTP+D2 is less pronounced than for ETN and ETN+D2. However, a reduction in the uncertainty region is still observed, along with an improved alignment between the mean and experimental RDFs. As in the case of density prediction, we associate the higher accuracy of the MLIP+D’s with reduced oscillations observed on the binding curves (Figure 5).
Overall, we observed that the CCl4 and CH4 RDFs were not sensitive to the incorporation of dispersion corrections in the functional form of the MLIPs. However, it had a significant effect on the RDF for toluene, highlighting that this effect depends on system complexity.
Additionally, we want to emphasize that the MLIP+D models still have advantages for the extended-cutoff MLIP(l)’s. In fact, employing a smaller $R_{\rm cut}$ while incorporating explicit dispersion corrections improves the computational efficiency of potential fitting and subsequent simulations. This efficiency stems from the reduced computational cost of DFT calculations for smaller configurations, as well as the faster training and evaluation of energies and forces with potentials that have a smaller cutoff radius.
It should be also noted that we started this study by calculating binding energies, density, and RDF using MLIPs with $R_{\rm cut}=5.5$ Å for the CCl4 system. We found that ETN+D3 outperformed ETN+D2 only for the CCl4 binding curves and, for this reason, we did not use ETN+D3 for other tests provided here.
Finally, the remaining deviations between the predicted densities and RDFs and experimental ones arises not only from MLIP fitting errors but also from intrinsic limitations of the PBE density functional, the dispersion correction schemes, and the neglect of many-body and quantum nuclear effects. While addressing these factors could potentially enhance the accuracy compared to experiment, it was beyond the scope of the present study.
IV Conclusions
In this work, we considered the problem of using MLIPs to model systems with strong long-range dispersion interactions. A typical MLIP cutoff radius is 5–6 Å which is much less than the range of dispersion interactions, hence the idea to include explicit dispersion interactions on top of a general-purpose MLIP.
We studied the impact of explicit incorporation of dispersion interactions in MTPs and ETN potentials via the D2 and D3 correction schemes. The impact was measured based on training and validation errors, errors in predicting binding curves, density, and RDFs of a liquid phase of three systems – carbon tetrachloride (CCl4), methane (CH4), and toluene.
Our first finding is that the accuracy of a typical MLIP (with a cutoff radius of 5.5 Å) is significantly improved after incorporating explicit dispersion. We observe an improvement in accuracy for predicting energy/forces and binding curves, and this improvement consistently translates into a higher accuracy on the finite-temperature properties.
Comparing the D2 and D3 corrections, we found that the incorporation of D2 is sufficient, as it provides the same order of accuracy as D3 in the systems we tested. This finding suggests that D2 can be a practical choice for many applications, as D3 is more computationally expensive (or at least harder to implement efficiently).
Another finding is that for the two relatively simple systems, CCl4 and CH4, the improvement in accuracy can instead be achieved by simply increasing the cutoff radius of the local MTP and ETN to 7.5 Å, except for the full-width at the half-minimum metrics in CH4, in which the MLIP+D’s are still more accurate. We note that we did not systematically study the effect of increasing the cutoff radius beyond 7.5 Å (due to the fast increase in the cost of collecting the appropriate training set). For the toluene system, which has a sufficiently complex geometry, a simple increase in the cutoff radius was not enough to achieve the accuracy of the MLIP+D models, especially for the ETN model.
Supplementary material
See the supplementary material for additional data and supporting discussion of statements made in the main text.
Acknowledgements.
This work was in part supported by the Russian Science Foundation (grant number 23-13-00332, https://rscf.ru/project/23-13-00332/). Max Hodapp acknowledges the financial support under the scope of the COMET program within the K2 Center “Integrated Computational Material, Process and Product Engineering (IC-MPPE)” (Project No 886385); this program is supported by the Austrian Federal Ministries for Climate Action, Environment, Energy, Mobility, Innovation and Technology (BMK) and for Labour and Economy (BMAW), represented by the Austrian Research Promotion Agency (FFG), and the federal states of Styria, Upper Austria and Tyrol.
Author declarations
Conflict of interest
The authors have no conflicts to disclose.
Author Contributions
Olga Chalykh: Conceptualization (equal); Methodology (equal); Software (equal); Formal analysis; Writing - original draft; Writing - review & editing (supporting). Dmitry Korogod: Software (equal); Writing - review & editing (supporting). Ivan S. Novikov: Methodology (equal); Software (supporting); Writing - original draft (supporting); Writing - review & editing (equal). Max Hodapp: Methodology (equal); Software (supporting); Writing - original draft (supporting); Writing - review & editing (supporting). Nikita Rybin: Methodology (supporting); Writing - review & editing (supporting). Alexander V. Shapeev: Conceptualization (equal); Methodology (equal); Writing - review & editing (equal).
Data Availability Statement
The MLIP-2 code used for the fitting and simulations with MTP is publicly available in the GitLab repository at https://gitlab.com/ashapeev/mlip-2.
The MLIP-4 code used for the fitting and simulations with ETN is publicly available in the GitLab repository at https://gitlab.com/ashapeev/mlip-4.